### v2

# Chance-Constrained Covariance Steering for Discrete-Time Markov Jump Linear Systems

Abstract
In this paper, we propose a novel convex optimization framework to solve the optimal covariance steering problem for discrete-time Markov Jump Linear Systems (MJLS) with chance constraints. We derive the analytical expressions for the mean and covariance trajectories of time-varying discrete-time MJLS and show that they cannot be separated even without chance constraints, unlike the single-mode dynamics case. To solve the covariance steering problem, we propose a two-step convex optimization framework, which optimizes the mean and covariance subproblems sequentially. Further, we use Gaussian approximations to incorporate chance constraints and propose an iterative optimization framework to solve the chance-constrained covariance steering problem. Both problems are originally nonconvex, and we derive convex relaxations which are proved to be lossless at optimality using the Karush–Kuhn–Tucker (KKT) conditions. Numerical simulations demonstrate the proposed method by achieving target covariances while respecting chance constraints under Gaussian noise and Markovian jump dynamics.
I Introduction
Covariance steering is an important problem of designing a control policy that steers a system from an initial given distribution to a target distribution while minimizing a cost function. It was first proposed in [1] as an infinite horizon covariance control problem, and has attracted significant interest recently as a finite horizon covariance steering problem[2, 3, 4, 5].
The covariance steering formulation was previously applied to vehicle path planning [6] and spacecraft trajectory optimization[7, 8, 9]. Such applications typically require path constraints due to safety considerations and physical limitations and have previously been incorporated in the form of probabilistic chance constraints[10]. [11] proposed a lossless convex relaxation of the chance-constrained covariance steering problem, which was later shown [12] to outperform existing methods [4, 6] in computation time.
Despite recent remarkable advances in covariance steering theory for smooth dynamics, covariance steering for hybrid systems has been significantly underexplored. [13] studies covariance steering for a class of hybrid systems that involve state-triggered switching. However, such approaches cannot extend to tackle covariance steering for Markov jump linear systems (MJLS). MJLS is a subclass of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain. They have been extensively studied in the literature with applications in fault-tolerant control [14], network control [15], and macroeconomic models [16]. [17] is an excellent reference on the classical results for discrete-time MJLS.
Constrained control of MJLS has been an active area of research, e.g. with state and control polytopic constraints for linear time-invariant[18] and periodic systems[19], and with second-moment constraints[20]. Chance constraints have previously been incorporated for MJLS[21], but they only consider time-invariant systems in a receding horizon framework using probabilistic reachable sets, which might not scale well with higher dimensions or longer time horizons. Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints and boundary conditions in a computationally tractable manner. To the authors’ best knowledge, such formulations have previously not been investigated in the context of MJLS.
In this work, we focus on the optimal covariance steering problem for MJLS with chance constraints. This is motivated by the need for probabilistic path planning approaches which face abrupt changes in dynamics, such as spacecrafts[22] and robots[20] with sudden actuator failure. The contributions of this work are threefold: (1) Demonstrate the coupling of MJLS mean and covariance dynamics, showing why mean-covariance separation of single-mode systems does not extend to MJLS, (2) Derive a lossless convex relaxation to solve the MJLS covariance steering problem in a computationally efficient manner, and (3) Incorporate chance constraints in the covariance steering framework and propose a sequential convex programming approach to solve the problem.
Notation
[…]
II MJLS Dynamics
We consider a discrete-time MJLS defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a deterministic bias $c_{k}$ and a zero-mean Gaussian noise $w_{k}\sim\mathcal{N}(\mathbf{0},{I})$ given by:
$x_{k+1}=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+c_{k}+G_{k}(r_{k})w_{k}$ (1)
where $x\in\mathbb{R}^{n_{x}}$ and $u_{k}(i)\in\mathbb{R}^{n_{u}}$ are the state and control at time $k$ respectively. $r_{k}$ denotes the active mode of a homogeneous Markov chain taking values in $\Omega=\{1\ldots N\}$ with transition probability matrix $\mathbb{P}$ whose ($i$,$j$)th entry is $p_{ij}$. We define the following variables to study MJLS dynamics:
$\rho_{k}(i)\triangleq\mathbb{E}\left[\mathds{1}_{\{r_{k}=i\}}\right],$ $q_{k}(i)\triangleq\mathbb{E}\left[x_{k}\mathds{1}_{r_{k}=i}\right],$
$\mu_{k}\triangleq\mathbb{E}[x_{k}]=\sum_{i\in\Omega}q_{k}(i),$ $\mathcal{X}_{k}(i)\triangleq\mathbb{E}\left[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}\right],$
$X_{k}\triangleq\mathbb{E}[x_{k}x_{k}^{\top}]=\sum_{i\in\Omega}\mathcal{X}_{k}(i),$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
We assume that the initial probability distribution $\rho_{0}\in\mathbb{R}^{1\times N}$ is given. Therefore, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathbb{P}^{k}$.
Remark 1.
Note that $x_{k}$ is independent of $r_{k}$. Therefore, $q_{k}(i)=\mathbb{E}[x_{k}]\mathbb{E}[\mathds{1}_{r_{k}=i}]=\mu_{k}\rho_{k}(i)$, and $\mathcal{X}_{k}(i)=\rho_{k}(i)X_{k}$.
We assume a control policy of the following form:
$u_{k}(i)=\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k})$ (2)
Assumption 1.
$\rho_{k}(i)>0,\;\forall k\in\mathbb{Z}_{0}^{T},i\in\Omega$, i.e. the probability of any mode being active is non-zero at all times.
Assumption 2.
The covariance of the given MJLS is positive definite at all time steps, i.e. $\Sigma_{k}\succ 0$ $\forall k\in\mathbb{Z}_{0}^{T}$.
II-A Mean Propagation
We now focus on MJLS mean dynamics. We know:
$q_{k+1}$ $(j)=\mathbb{E}\left[x_{k+1}\mathds{1}_{\{r_{k+1}=j\}}\right]$
$=\sum_{i\in\Omega}\mathbb{E}\left[\left({A}_{k}(i)x_{k}+B_{k}(i)(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))\right.\right.$
$\left.\left.\hskip 85.35826pt+c_{k}+G_{k}(i)w_{k}\right)p_{ij}\mathds{1}_{\{r_{k}=i\}}\right]$
$=\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (3)
Using the definition of $\mu_{k}$, eq. 3 can also be written as:
$\mu_{k+1}=\sum_{j\in\Omega}\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (4)
We define a mean trajectory $\tau$ as a finite sequence of $(N+1)$-tuples $\{\mu_{k},\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in\mathbb{Z}_{0}^{T-1}}$ which satisfy eq. 4.
II-B Covariance Propagation
For covariance propagation, we first consider $\mathcal{X}_{k+1}(j)$:
$\mathcal{X}_{k+1}(j)=\mathbb{E}\left[x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k+1}=j}\right]$ (5)
$=\sum_{i\in\Omega}\left\{p_{ij}\mathbb{E}\left[\left(A_{k}(i)x_{k}+B_{k}(i)u_{k}(i)+c_{k}+G_{k}(i)w_{k}\right)\right.\right.$
$\left.\left.\left(A_{k}(i)x_{k}+B_{k}(i)u_{k}(i)+c_{k}+G_{k}(i)w_{k}\right)^{\top}\mathds{1}_{r_{k}=i}\right]\right\}$
Using identities derived in the appendix, we can simplify eq. 5, and substitute that in $\Sigma_{k+1}=\sum_{j\in\Omega}\mathcal{X}_{k+1}(j)-\mu_{k+1}\mu_{k+1}^{\top}$ to get:
${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)\right.$
$\left.+A_{k}(i)(\mu_{k}\overline{u}^{\top}_{k}(i)+\Sigma_{k}K_{k}^{\top}(i))B^{\top}_{k}(i)+A_{k}(i)\mu_{k}c^{\top}_{k}\right.$
$\left.+B_{k}(i)(\mu_{k}\overline{u}^{\top}_{k}(i)+\Sigma_{k}K^{\top}_{k}(i))^{\top}A_{k}^{\top}(i)+B_{k}(i)\overline{u}^{\top}_{k}(i)c^{\top}_{k}\right.$
$\left.+B_{k}(i)(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i)\Sigma_{k}K_{k}^{\top}(i))B_{k}^{\top}(i)+c_{k}c_{k}^{\top}\right.$
$\left.+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+G_{k}(i)G_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)\mu_{k}\mu_{k}^{\top}A^{\top}_{k}(i)\}\right]\Bigr{)}-\mu_{k+1}\mu_{k+1}^{\top}$ (6)
Remark 2.
Separately analyzing some terms inside the summation in eq. 6:
$A_{K}(i)\mu_{k}\mu_{k}^{\top}A_{K}^{\top}(i)+A_{k}(i)\mu_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+A_{k}(i)\mu_{k}c_{k}^{\top}$
$+B_{k}(i)\overline{u}_{k}(i)\mu^{\top}_{k}A_{k}^{\top}(i)+B_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)$
$+B_{k}(i)\overline{u}_{k}(i)c^{\top}_{k}+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}\overline{u}^{\top}_{k}B_{k}^{\top}(i)+c_{k}c_{k}^{\top}$
$=(A_{k}(i)\mu_{k}+B_{k}\overline{u}_{k}(i)+c_{k})(A_{k}(i)\mu_{k}+B_{k}\overline{u}_{k}(i)+c_{k})^{\top}$
Note these terms are linear in $p_{ij}$ while $\mu_{k+1}\mu_{k+1}^{\top}$ involves the product $p_{ij}p_{ij}$, explaining why they don’t cancel out. This demonstrates why the mean-covariance separation of single-mode dynamics (shown in [10]) does not extend to MJLS.
III Unconstrained MJLS Covariance Steering
III-A Optimal Control Problem
The unconstrained optimal control problem is given by:
$\min$ $\;J=\sum_{k=0}^{T-1}J_{k}$ (7)
$s.t$$.\;\;\;\forall k\in\mathbb{Z}_{0}^{T-1},$
$x_{k+1}$ $=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+G_{k}(r_{k})c_{k}$
$x_{0}$ $\sim\mathcal{N}(\overline{\mu}_{0},\overline{\Sigma}_{0})$
$\mathbb{E}[x_{T}]=$ $\mu_{f},\;\Sigma_{T}\preceq\Sigma_{f}$
where $J_{k}=\mathbb{E}\left[x_{k}Q_{k}x_{k}^{\top}+u_{k}(i)R_{k}u_{k}^{\top}(i)\right]$ with $Q_{k}\succeq 0$ and $R_{k}\succ 0$, $\forall k\in\mathbb{Z}_{0}^{T-1}$. Simplifying $J_{k}$, we get:
$J_{k}$ $=\sum_{i\in\Omega}\mathbb{E}\left[(x_{k}^{\top}Q_{k}x_{k}+u^{\top}_{k}(i)R_{k}u_{k}(i))\mathds{1}_{r_{k}=i}\right]$
$=\sum_{i\in\Omega}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
Using the identities derived in appendix A, we can rewrite $J_{k}$ as:
$J_{k}$ $=\sum_{i\in\Omega}\rho_{k}(i)\left[\mathrm{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k})+\mathrm{Tr}(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right.$ (8)
$\left.\hskip 28.45274pt+\mathrm{Tr}({\Sigma}_{k}Q_{k})+\mathrm{Tr}(K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)R_{k})\right]$
III-B Convex Relaxation
The nonlinear terms in eq. 6 result in a non-convex optimization landscape for eq. 7. To solve it in a computationally tractable manner, we propose a two-step optimization process, which optimizes the mean and covariance variables sequentially using convex relaxations. We first introduce the following change of variables to convexify eq. 6:
$L_{k}(i)$ $=K_{k}(i){\Sigma}_{k},\;$ $Y_{k}(i)$ $=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$ (9)
Rearranging eq. 6 using the above-mentioned terms,
${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)+G_{k}(i)G_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)L^{\top}_{k}(i)B_{k}^{\top}(i)+B_{k}(i)L_{k}(i)A_{k}^{\top}(i)+B_{k}(i)Y_{k}(i)B_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)\mu_{k}\overline{u}^{\top}_{k}(i)B^{\top}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)\mu^{\top}_{k}A_{k}^{\top}(i)+A_{k}(i)\mu_{k}c_{k}^{\top}\right.$
$\left.+B_{k}(i)\overline{u}_{k}(i)c^{\top}_{k}+c_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+B_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)\right.$
$\left.+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}c_{k}^{\top}+A_{k}(i)\mu_{k}\mu_{k}^{\top}A^{\top}_{k}(i)]\}\right]\Bigr{)}-\mu_{k+1}\mu_{k+1}^{\top}$
For a given mean trajectory $\tau$, we define covariance trajectory $\Xi$ as a finite sequence of $(2N+1)$-tuples $\{\Sigma_{k},L_{k}(1)\ldots L_{k}(N),Y_{k}(1)\ldots Y_{k}(N)\}_{k\in\mathbb{Z}_{0}^{T-1}}$ which satisfy section III-B. Thus, section III-B can be expressed as
$\Sigma_{k+1}=F_{\tau}(\Xi)$ (11)
where $F_{\tau}(\Xi)$ is a linear operator in $\Xi$ for a given $\tau$. We relax the non-convex definition of $Y_{k}(i)$ in eq. 9 as:
$C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)\preceq 0$ (12)
which can be recast as a LMI using Schur’s complement:
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}\succeq 0$ (13)
We use these new variables in the objective function as follows:
$J$ $=J(\tau)+J(\Xi)$ (14)
where
$J(\tau)=\sum_{k=0}^{T-1}$ $\sum_{i\in\Omega}\rho_{k}(i)\left[\text{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k}+\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right]$ (15a)
$J(\Xi)$ $=\sum_{k=0}^{T-1}\sum_{i\in\Omega}\rho_{k}(i)\left[\text{Tr}({\Sigma}_{k}Q_{k}+Y_{k}(i)R_{k})\right]$ (15b)
III-C Two-step Convex Optimization Framework
We now define the two subproblems which are solved sequentially. Therefore, the two-step optimization process first solves Problem 1 and then uses its solution to solve Problem 2.
Problem 1.
Solve:
$\min_{\mu_{k},\overline{u}_{k}(i)}J(\tau)$ (16a)
[…]
$\mu_{0}=\overline{\mu}_{0},\;\mu_{T}=\mu_{F}$ (16c)
Problem 2.
Given a mean trajectory $\tau$, solve:
$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i)}J(\Xi)$ (17a)
[…]
$H_{K}^{\Sigma}=\Sigma_{k+1}$ $-F_{\tau}(\Xi)=0$ (17b)
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}$ $\succeq 0$ (17c)
$\Sigma_{0}=\overline{\Sigma}_{0},$ $\quad\Sigma_{T}\preceq\Sigma_{f}$ (17d)
Assumption 3.
Slater’s condition is satisfied for Problem 2, i.e. there exists a strictly feasible solution for Problem 2.
Lemma 1 (Lemma 1 in [11]).
Let $A$ and $B$ be $n\times n$ symmetric matrices with $A\succeq 0,B\preceq 0$, and $\mathrm{Tr}(AB)=0$. If $B$ has at least one nonzero eigenvalue, then, $A$ is singular.
Theorem 1.
Under Assumptions 1, 2 and 3, the optimal solution to Problem 2 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$, hence, lossless.
Proof.
For Problem 2, we can define the Lagrangian as:
$\mathcal{L}(\cdot)=J(\Xi)+$ $\sum_{k=0}^{T-1}\left[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))\right\}\right]$ (18)
$+\mathrm{Tr}(\lambda^{\top}_{F}(\Sigma_{T}-\Sigma_{f}))$
where $C_{k}(i)$ is defined in eq. 12, $\Lambda_{k}$, $M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively. $\Lambda_{k}$ is symmetric because of symmetry in eq. 17b which shares the same symmetry properties with eqs. 11 and III-B. $M_{k}$ and $\lambda_{F}$ are symmetric by definition. The relevant first-order KKT conditions for optimality are:
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}$ $=2\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}=0$ (19a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}$ $=\rho_{k}(i)R_{k}-M_{k}(i)+\sum_{j\in\Omega}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (19b)
$\mathrm{Tr}($ $M_{k}^{\top}(i)C_{k}(i))=0,\quad C_{k}\preceq 0,\quad M_{k}\succeq 0$ (19c)
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies the corresponding Lagrange multiplier $M_{k}(i)$ is singular from Lemma 1 and eq. 19c. Rearranging eq. 19a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$. Using this and the substitutions defined earlier in eq. 9, we simplify eq. 19b as: $\rho_{k}(i)R_{k}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Taking the determinant of both sides, we get the following:
$\mathrm{det}(\rho_{k}(i)R_{k})=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
which contradicts $\rho_{k}(i)R_{k}\succ 0$ from Assumption 1. Therefore, we have $C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)=0,\;\forall k,i.$
Remark 3.
If we attempt to carry out similar relaxations for the mean quadratic terms to relax the covariance propagation equation (6) without assuming the mean trajectory $\tau$ is given, we find that it is not possible to prove that these relaxations are lossless. In fact, we get a condition where $M_{k}^{(3)}=-M^{(4)}_{k}$, where $M_{k}^{(3)}$ and $M^{(4)}_{k}$ are Lagrange multipliers for respective relaxations. Since they are both positive semi-definite, it implies $M_{k}^{(3)}=M^{(4)}_{k}=0$. Thus, we cannot use our existing arguments to prove the relaxation is lossless. Numerical simulations also demonstrated this lossy behavior, motivating our two-step sequential mean-covariance optimization approach.
IV Chance-Constrained MJLS Covariance Steering
We now consider chance constraints for both state and control. For state, we consider intersection of hyperplane constraints and tube-like constraints around the mean trajectory. Mathematically, they’re expressed as:
$P(a^{\top}_{j}x_{k}+b_{j}\leq 0,\forall j)$ $\geq 1-\delta_{x}$ (20a)
$P(\|x_{k}-\mu_{k}\|\leq d_{\max})$ $\geq 1-\epsilon_{x}$ (20b)
Likewise, for control, we consider
$P(p^{\top}_{j}u_{k}(i)+q_{j}\leq 0,\forall j)$ $\geq 1-\delta_{u}(i)$ (21a)
$P(\|u_{k}(i)\|\leq u_{\max}(i))$ $\geq 1-\epsilon_{u}(i)$ (21b)
Previous works [23, 8] have focused on deterministic and convex reformulations of eqs. 20 and 21, but they usually require Gaussian dynamics. MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) In this work, we approximate the MJLS state by a single Gaussian distribution. To achieve this, we seek to minimize the Kullback-Leibler divergence [24] between the MJLS distribution and the approximating Gaussian. It has been shown previously [25] that this is achieved by performing moment matching, i.e. setting the first two moments of the approximating Gaussian to be the same as the original distribution which is being approximated. We leverage this useful property to approximate the state and control of MJLS using Gaussian distributions. We approximate $x_{k}$ by $\mathcal{N}(\mu_{k},\Sigma_{k})$, using moment matching to the approximating Gaussian. Similarly, for control in mode $i$, $\mathbb{E}[u_{k}(i)]=\overline{u}_{k}(i)$ and $\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}]=K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)=Y_{k}(i)$, therefore $u_{k}(i)$ can be approximated by $\mathcal{N}(\overline{u}_{k}(i),Y_{k}(i))$. We now discuss the deterministic formulation of eqs. 20 and 21.
IV-A Deterministic Formulation
Lemma 2 (Lemma 2 in [8]).
Let ${\xi}\sim\mathcal{N}(\bar{{\xi}},P_{\xi})\in\mathbb{R}^{n_{\xi}}$ and $\epsilon\in(0,0.5)$. Then, $\mathbb{P}[a_{j}^{\top}{\xi}+b_{j}\leq 0,\;\forall j]\geq 1-\epsilon$ is implied by
$a_{j}^{\top}\bar{{\xi}}+b_{j}+m_{\mathcal{N}}(\epsilon_{j})\sqrt{a_{j}^{\top}P_{\xi}a_{j}}\leq 0,\;\forall j,\quad\sum_{j}\epsilon_{j}\leq\epsilon$ (22)
where $m_{\mathcal{N}}(\epsilon)=Q_{X\sim\mathcal{N}}(1-\epsilon)$ denotes the quantile function of the standard normal distribution at probability $1-\epsilon$.
Lemma 3 (Lemma 3 in [8]).
Let ${\xi}\sim\mathcal{N}(\bar{{\xi}},P_{\xi})\in\mathbb{R}^{n_{\xi}}$, $\epsilon\in(0,1)$, and $\gamma>0$. Then, $\mathbb{P}[\|{\xi}\|\leq\gamma]\geq 1-\epsilon$ is implied by
$\|\bar{{\xi}}\|+m_{\chi^{2}}(\epsilon,n_{\xi})\sqrt{\lambda_{\max}(P_{\xi})}\leq\gamma$ (23)
$m_{\chi^{2}}(\epsilon,n_{\xi})=\sqrt{Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)}$. Here, $Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)$ denotes the quantile function of the chi-squared distribution with $n_{\xi}$ degrees of freedom, evaluated at probability $(1-\epsilon)$.
Note that while eqs. 22 and 23 are convex in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
$m^{2}_{\mathcal{N}}(\epsilon_{j}){a^{\top}_{j}P_{\xi}a_{j}}\leq(a_{j}^{\top}\bar{\xi}+b_{j})^{2}$ (24)
$m^{2}_{\chi^{2}}(\epsilon,n_{\xi}){\lambda_{\max}(P_{\xi})}\leq(\gamma-\|\bar{\xi}\|)^{2}$ (25)
which are convex in the covariance $P_{\xi}$.
IV-B Iterative Convex Control Synthesis Framework
Using the deterministic formulation of chance constraints in Lemmas 2 and 3, we can now incorporate them in our two-step optimization framework. Note that the chance constraints introduce two-way coupling in the problem, while previously only the covariance problem depended on the mean trajectory and not vice-versa. To solve this, we propose a hierarchical optimization framework which iteratively solves the mean and covariance optimization problem by relaxing the chance constraints. We relax the constraints by adding a slack variable, and then penalize its 1-norm in the objective function. Using the 1-norm promotes sparsity of the slack variable, and has demonstrated favorable convergence properties in other hierarchical optimization frameworks [26, 27]. We now define the relaxed mean and covariance problems as follows:
Problem 3.
Given a covariance trajectory $\Xi$, solve:
$\min_{\mu_{k},\overline{u}_{k}(i),\beta_{m}}J_{\mu}^{c}$ (26a)
[…]
$a_{j}^{\top}\mu_{k}$ $+b_{j}+m_{\mathcal{N}}(\delta_{x,j})\sqrt{a_{j}^{\top}\Sigma_{k}a_{j}}\leq\beta_{1,k}$ (26c)
$p_{j}^{\top}\overline{u}_{k}(i)$ $+q_{j}+m_{\mathcal{N}}(\delta_{u,j}(i))\sqrt{p_{j}^{\top}Y_{k}(i)p_{j}}\leq\beta_{2,k}$ (26d)
$\|\overline{u}_{k}(i)\|+$ $m_{\chi^{2}}(\epsilon_{u}(i),n_{u})\sqrt{\lambda_{\max}(Y_{k}(i))}\leq u_{\max}(i)+\beta_{3,k}$ (26e)
where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $J_{\mu}^{c}=J(\tau)+\sum_{m}\alpha_{m}\|\beta_{m}\|_{1}$, with $J(\tau)$ defined in eq. 15a. $\beta_{m,k}>0$ is the slack variable for $m$-th constraint at time step $k$ and $\beta_{m}=[\beta_{m,0}\ldots\beta_{m,T-1}]^{\top}$. $\alpha_{m}>0$ is the weight corresponding to the slack variable vector 1-norm $\|\beta_{m}\|_{1}$ in $J_{\mu}^{c}$.
Problem 4.
Given a mean trajectory $\tau$, solve:
$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i),\zeta_{m}}J^{c}_{\Sigma}\hskip 56.9055pt$ (27a)
[…]
$H_{K}^{\Sigma}=\Sigma_{k+1}-F_{\tau}(\Xi)$ $=0$ (27b)
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}$ $\succeq 0$ (27c)
$\Sigma_{0}=\overline{\Sigma}_{0},$ $\quad\Sigma_{T}\preceq\Sigma_{f}$ (27d)
$m^{2}_{\mathcal{N}}(\delta_{x,j}){a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $\leq\zeta_{1,j},$ (27e)
$m^{2}_{\mathcal{N}}(\delta_{u,j}){p_{j}^{\top}Y_{k}(i)p_{j}}-(p_{j}^{\top}\mu_{k}+d_{j})^{2}$ $\leq\zeta_{2,j},$ (27f)
$m^{2}_{\chi^{2}}(\epsilon_{x},n_{x})\lambda_{\max}(\Sigma_{k})$ $\leq\zeta_{3,j}$ (27g)
$m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $\leq\zeta_{4,j}$ (27h)
where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $u_{m}=u_{\max}-\|\overline{u}_{k}(i)\|$, and $J_{\Sigma}^{c}=J(\Xi)+\sum_{j}\alpha_{m}\|\zeta_{m}\|_{1}$ with $J(\Xi)$ defined in eq. 15b. $\zeta_{m,k}>0$ is the slack variable for $m$-th constraint at time $k$ and $\zeta_{m}=[\zeta_{m,0}\ldots\zeta_{m,T-1}]^{\top}$. $\alpha_{m}>0$ is the weight for slack variable vector 1-norm $\|\zeta_{m}\|_{1}$ in $J_{\Sigma}^{c}$.
Assumption 4.
Slater’s condition is satisfied in Problem 4 for $\zeta_{m}=0$. This ensures the convex relaxation in eq. 27c remains lossless while satisfying chance constraints at $\zeta_{m}=0$.
We now show that the convex relaxation $Y_{k}(i)\succeq L_{k}(i)\Sigma_{k}^{-1}L_{k}^{\top}(i)$ remains lossless in Problem 4.
Theorem 2.
Under Assumptions 1, 2 and 4, the optimal solution to Problem 4 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$.
Proof.
We can define the Lagrangian for Problem 4 as follows:
$\mathcal{L}(\cdot)=$ $J_{\Sigma}^{c}+\sum_{k=0}^{T-1}\Bigl{[}\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\lambda_{x,n}e_{x,n}+\lambda_{x,h}e_{x,h}$ (28)
$+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))+\lambda_{u,n}(i)e_{u,n}(i)\right.$
$\left.+\lambda_{u,h}(i)e_{u,h}(i)\right\}\Bigr{]}+\mathrm{Tr}(\lambda^{\top}_{F}(\Sigma_{T}-\Sigma_{f}))$
where $C_{k}(i)$ is given in eq. 12 and $e_{x,h}$, $e_{u,h}$, $e_{x,n}$, and $e_{u,n}$ represent eqs. 27e, 27f, 27g and 27h respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers. $\Lambda_{k}$,$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1. The relevant first-order KKT conditions for optimality are:
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}=2\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}=0$ (29a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}=\rho_{k}(i)R_{k}-M_{k}(i)+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})v_{k}(i)v_{k}^{\top}(i)$
$+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}+\sum_{j\in\Omega}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (29b)
$\mathrm{Tr}(M_{k}^{\top}(i)C_{k}(i))=0,\quad C_{k}\preceq 0,\quad M_{k}\succeq 0,$
$\hskip 56.9055pt\lambda_{x,n},\lambda_{x,h},\lambda_{u,h},\lambda_{u,n}\geq 0$ (29c)
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$ and $v_{k}(i)$ is the eigenvector of $Y_{k}(i)$ corresponding to its maximum eigenvalue. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies $M_{k}(i)$ is singular from Lemma 1 and eq. 29c. Rearranging eq. 29a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$. Using this and the substitutions defined in eq. 9, we simplify eq. 29b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Representing the left hand side as $\Delta_{k}(i)$ and taking the determinant of both sides, we get:
$\mathrm{det}(\Delta_{k}(i))=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
From Assumption 1, we have that $\rho_{k}(i)R_{k}\succ 0$. Now we know $\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}\geq 0$ which implies $\Delta_{k}(i)>0$, resulting in a contradiction. Hence, $C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)=0\;\forall k\in\mathbb{Z}_{0}^{T-1},i\in\Omega$. ∎
IV-C Sequential Convex Optimization Framework
We now describe the sequential convex optimization framework which is used to design optimal control laws for MJLS covariance steering with chance constraints. The framework requires a mean trajectory to be initialized (which is obtained by optimizing the unconstrained mean trajectory in Problem 1) and then solve Problem 4 and Problem 3 iteratively with increasing weights on the slack variables until they reach negligible values. This framework is described in Algorithm 1.
Algorithm 1 MJLS Chance-constrained covariance steering
1:Set tol, initial weights $\{\alpha_{m}\}$, and $\eta>1$;
2:Initialize mean trajectory $\tau$ by solving Problem 1;
3:while NotConverged do
4:     Solve Problem 4 with last stored mean trajectory $\tau$ to get new covariance trajectory $\Xi$;
5:     Compute new mean trajectory $\tau$ by solving Problem 3 using revised covariance trajectory $\Xi$ from previous step ;
6:     Store values of slack variables vector $\beta_{m}$ and $\zeta_{m}$;
7:     if $\max_{m}\|\beta_{m}\|_{\infty}$$>$tol or $\max_{m}\|\zeta_{m}\|_{\infty}$ $>$tol then
8:         increase weights $\{\alpha_{m}\}$ by a factor $\eta$;
9:     else
10:         Solution converged. break;
11:return $\tau$ and $\Xi$
V Numerical Simulations
We use the following dynamics:
$A_{k}(1)$ $=\begin{bmatrix}-0.2&1\\ -0.1&0.1\end{bmatrix}$ $B_{k}(1)$ $=\begin{bmatrix}1&0\\ 2&0\end{bmatrix}$ $G_{k}(1)$ $=\begin{bmatrix}0.1&0\\ 0&0.1\end{bmatrix}$
$A_{k}(2)$ $=\begin{bmatrix}0.2&0.1\\ -0.5&0.1\end{bmatrix}$ $B_{k}(2)$ $=\begin{bmatrix}0&1\\ -1&2\end{bmatrix}$ $G_{k}(2)$ $=\begin{bmatrix}0.5&0\\ 0&0.5\end{bmatrix}$
$\mathbb{P}$ $=\begin{bmatrix}0.7&0.3\\ 0.9&0.1\end{bmatrix}$ $c_{k}$ $=\begin{bmatrix}0.01\\ 0.01\end{bmatrix}$ $T$ $=6$
The initial and final conditions are set as $\rho_{0}=[0.5,0.5]$, $\overline{\mu}_{0}=[40,10]^{\top}$, $\overline{\Sigma}_{0}=6I$ and $\mu_{f}=[5,10]^{\top}$, $\Sigma_{f}=2.5I$. We set a half-space state chance constraint of $P(a^{\top}x>b)\geq 0.95$, where $a=[1,0]^{\top}$ and $b=-10$ and a control norm bound constraint $P(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=5$. We set the cost weight matrices $Q_{k}=0$ and $R_{k}=I$ $\forall k\in\mathbb{Z}_{0}^{T-1}$.
We use the CVXPY package[28] with MOSEK solver[29] to run Algorithm 1 on a 2023 Macbook Pro M1 Pro laptop with 16 GB of RAM. All initial weights are set at $\alpha_{m}=1e2$, with $\eta=5$ and $tol=1e-6$. It takes 3.43 seconds and $3$ iterations for Algorithm 1 to converge to the preset tolerance values. Figures 1(a) and 1(b) plot 1000 Monte Carlo trajectories for this problem without and with state chance constraints, respectively. Both figures use the control norm chance constraint. The initial covariance, terminal covariance, and the mean trajectory are shown in green, red, and blue, respectively, while Monte Carlo trajectories are shown in grey. The unshaded area indicates the feasible region under chance constraints, enforced using the state half-space constraint. Figure 2 shows the control norm history for the Monte Carlo simulations. None of the samples disobeyed the state chance constraints, while the control norm exceeded $u_{\max}$ in $1.3\%$ of the total sample size, which meets the prescribed risk bound of $5\%$. These numerical results demonstrate the performance of the controller to steer the covariance within the final target value in the given time horizon, while respecting the chance constraints. Moreover, the state trajectories in Figure 1 demonstrate the multi-modal nature of the problem because of the Markovian jump dynamics and mode-dependent controllers. We also note the number of state distribution modes increasing over time, most apparent in Figure 1(a), clearly demonstrating the exponential Gaussian mixture distribution of MJLS, which is hard to control using conventional stochastic optimal control approaches.
(a) unconstrained case
(b) with chance constraints
Figure 1: State trajectories for Monte Carlo simulations with the state chance constraint $P(x_{1}\geq-10)\geq 0.95$ (shaded region: infeasible.) Both cases use $P(\|u_{k}(i)\|\leq 5)\leq 0.95$.
Figure 2: Control Norm History for the Monte Carlo simulations with state chance constraint $P(x_{1}\geq-10)\geq 0.95$.
VI Conclusion
In this paper, we consider the optimal covariance steering problem for MJLS under chance constraints. We demonstrate the coupled nature of the problem and decompose the problem into a mean-covariance problem, which can be solved in a two-step convex optimization framework. We propose lossless convex relaxations of the covariance subproblem in both unconstrained and chance-constrained cases and discuss an iterative convex optimization framework to solve the chance-constrained covariance steering problem. Numerical simulations for the covariance steering framework are presented.
Appendix A MJLS identities
We show some of the identities which are used in deriving the covariance propagation equation for MJLS.
$\mathbb{E}[x_{k}u_{k}^{\top}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[x_{k}(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))^{\top}\mathds{1}_{r_{k}=i}]$
$=\rho_{k}(i)\left[\mu_{k}\overline{u}^{\top}_{k}(i)+{\Sigma}_{k}K^{\top}_{k}(i)\right]$ (30)
$\mathbb{E}[u_{k}(i)c_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))c_{k}^{\top}\mathds{1}_{r_{k}=i}]$
$=\overline{u}_{k}(i)c_{k}^{\top}\mathbb{E}[\mathds{1}_{r_{k}=i}]+0=\overline{u}_{k}(i)c_{k}^{\top}\rho_{k}(i)$ (31)
$\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))(\overline{u}_{k}(i)+$
$\hskip 142.26378ptK_{k}(i)(x_{k}-\mu_{k}))^{\top}\mathds{1}_{r_{k}=i}]$
$=\mathbb{E}[\mathds{1}_{r_{k}=i}]\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+\overline{u}_{k}(i)(\mathbb{E}[x^{\top}_{k}\mathds{1}_{r_{k}=i}]-$
$\mu^{\top}_{k}\mathbb{E}[\mathds{1}_{r_{k}=i}])K_{k}^{\top}(i)+K_{k}(i)(\mathbb{E}[x_{k}\mathds{1}_{r_{k}=i}]-\mu_{k}\mathbb{E}[\mathds{1}_{r_{k}=i}])\overline{u}_{k}^{\top}(i)$
$\hskip 42.67912pt+K_{k}(i)\mathbb{E}[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
$=\rho_{k}(i)\left[\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i){\Sigma}_{k}K_{k}^{\top}(i)\right]$ (32)