### v3

# Chance-Constrained Covariance Steering for Discrete-Time Markov Jump Linear Systems

Abstract
In this paper, we propose a novel convex optimization framework to solve the optimal covariance steering problem for discrete-time Markov Jump Linear Systems (MJLS) with chance constraints. We derive the analytical expressions for the mean and covariance trajectories of time-varying discrete-time MJLS and show that they cannot be separated even without chance constraints, unlike the single-mode dynamics case. To solve the covariance steering problem, we propose a two-step convex optimization framework, which optimizes the mean and covariance subproblems sequentially. Further, we use Gaussian approximations to incorporate chance constraints and propose an iterative optimization framework to solve the chance-constrained covariance steering problem. Both problems are originally nonconvex, and we derive convex relaxations which are proved to be lossless at optimality using the Karush–Kuhn–Tucker (KKT) conditions. Numerical simulations demonstrate the proposed method by achieving target covariances while respecting chance constraints under Gaussian noise and Markovian jump dynamics.
I Introduction
Covariance steering is an important problem of designing a control policy that steers a system from an initial distribution to a target distribution while minimizing a cost function. It was first proposed in [1] as an infinite horizon problem and has attracted significant interest recently[2, 3, 4, 5].
The covariance steering formulation was previously applied to vehicle path planning [6] and spacecraft trajectory optimization[7, 8, 9]. Such applications typically require path constraints due to safety considerations and physical limitations and have previously been incorporated in the form of probabilistic chance constraints[10]. [11] proposed a lossless convex relaxation of the chance-constrained covariance steering problem, which was later shown [12] to outperform existing methods [4, 6] in computation time.
Despite recent remarkable advances in covariance steering theory for smooth dynamics, covariance steering for hybrid systems has been significantly underexplored. [13] studies covariance steering for a subclass of hybrid systems but don’t extend to Markov jump linear systems (MJLS). MJLS is a type of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain. They have been extensively studied in the literature[14] with applications in fault-tolerant control [15], networks [16], and macroeconomic models [17].
Constrained control of MJLS has been an active area of research, and has benefited from the enormous progress in constrained control of single mode systems [bemporad_explicit_2002, 18] e.g. with state and control polytopic constraints for linear time-invariant[19] and periodic systems[20], and with second-moment constraints[21]. Chance constraints have previously been incorporated for MJLS[22], but they only consider time-invariant systems in a receding horizon framework using probabilistic reachable sets, which might not scale well with higher dimensions or longer time horizons. Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints in a computationally tractable manner. To the authors’ best knowledge, such formulations have previously not been investigated in the context of MJLS.
In this work, we focus on the optimal covariance steering problem for MJLS with chance constraints. This is motivated by the need for probabilistic path planning approaches which face abrupt changes in dynamics, such as spacecrafts[23] and robots[21] with sudden actuator failure. The contributions of this work are threefold: (1) Demonstrate the coupling of MJLS mean and covariance dynamics, showing why mean-covariance separation of single-mode systems does not extend to MJLS, (2) Derive a lossless convex relaxation to solve the MJLS covariance steering problem in a computationally efficient manner, and (3) Incorporate chance constraints in the covariance steering framework and propose a sequential convex programming approach to solve the problem.
Notation
[…]
II Problem Statement
We consider a discrete-time MJLS defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a deterministic bias $c_{k}$ and a zero-mean Gaussian noise $w_{k}\sim\mathcal{N}(\mathbf{0},{I})$ given by:
$x_{k+1}=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+c_{k}+G_{k}(r_{k})w_{k}$ (1)
where $x\in\mathbb{R}^{n_{x}}$ and $u_{k}(i)\in\mathbb{R}^{n_{u}}$ are the state and control at time $k$ respectively. $r_{k}$ denotes the active mode of a homogeneous Markov chain taking values in $\Omega=\{1\ldots N\}$ with transition probability matrix $\mathbb{P}$ whose ($i$,$j$)th entry is $p_{ij}$. We define the following variables to study MJLS dynamics:
$\rho_{k}(i)\triangleq\mathbb{E}\left[\mathds{1}_{\{r_{k}=i\}}\right],$ $q_{k}(i)\triangleq\mathbb{E}\left[x_{k}\mathds{1}_{r_{k}=i}\right],$
$\mu_{k}\triangleq\mathbb{E}[x_{k}]=\sum_{i\in\Omega}q_{k}(i),$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
The mode probability distribution vector is given by $\rho_{k}=[\rho_{k}(1)\ldots\rho_{k}(N)]$. For a given initial mode distribution $\rho_{0}$, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathbb{P}^{k}$. We approximate $q_{k}(i)$ by $\rho_{k}(i)\mu_{k}$, which was found to be a reasonable approximation numerically. Similarly, we approximate $\mathbb{E}[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\mathds{1}_{r_{k}=i}]$ by $\rho_{k}(i)\Sigma_{k}$. The appendix section elaborates on this approximation and the consequent justification. We now define the MJLS covariance steering optimal control problem below:
Problem 1.
Solve:
$\min_{x_{k},u_{k}(i)}$ $\;J=\sum_{k=0}^{T-1}\mathbb{E}\left[x_{k}Q_{k}x_{k}^{\top}+u_{k}(i)R_{k}u_{k}^{\top}(i)\right]$ (2a)
$s.t$$.\;\;\;\forall k\in\mathbb{Z}_{0:T-1},$
$x_{k+1}$ $=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+c_{k}+G_{k}(r_{k})w_{k}$ (2b)
$x_{0}\sim\mathcal{N}(\overline{\mu}_{0},$ $\overline{\Sigma}_{0}),\;\mathbb{E}[x_{T}]=\mu_{f},\;\Sigma_{T}\preceq\Sigma_{f}$ (2c)
$P(x_{k}\in\mathcal{X})$ $\geq 1-\epsilon_{x},\quad P(u_{k}(i)\in\mathcal{U})\geq 1-\epsilon_{u}$ (2d)
where $Q_{k}\succeq 0$, $R_{k}\succ 0$ and $A_{k}(i)$ is invertible $\forall k\in\mathbb{Z}_{0:T},i\in\Omega$. The distribution $\rho_{0}$, $w_{k}$ and $x_{0}$ are independent.
Assumption 1.
$\rho_{k}(i)>0,\;\forall k\in{\mathbb{Z}_{0:T}},i\in\Omega$, i.e. the probability of any mode being active is non-zero at all times.
Assumption 2.
$\Sigma_{k}\succ 0$ $\forall k\in{\mathbb{Z}_{0:T}}$.
Problem 1 is hard to solve because of stochastic decision variables. To solve it in a computationally tractable manner, we focus on the mean and covariance propagation of the MJLS. We consider a mode-dependent affine feedback control policy:
${u_{k}(i)=\overline{u}_{k}(i)+K_{k}(i)\left(x_{k}-\overline{x}_{k}(i)\right)}$ (3)
Such control laws with a linear feedback law is common in covariance steering literature [11, 12] as it allows us to steer both the mean and covariance simultaneously. Focusing on the mean propagation first, we consider $q_{k}(i)$ below:
$q_{k+1}$ $(j)=\sum_{i\in\Omega}\mathbb{E}\left[\left({A}_{k}(i)x_{k}+B_{k}(i)(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))\right.\right.$
$\left.\left.\hskip 85.35826pt+c_{k}+G_{k}(i)w_{k}\right)p_{ij}\mathds{1}_{\{r_{k}=i\}}\right]$
$=\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (4)
Using the definition of $\mu_{k}$, eq. 4 can also be written as:
$\mu_{k+1}=\sum_{j\in\Omega}\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (5)
Similarly, covariance propagation is studied below:
$\Sigma_{k+1}$ $=\left(\sum_{i\in\Omega}\mathbb{E}[(x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k}=i}]\right)-\mu_{k+1}\mu_{k+1}^{\top}$ (6)
Substituting the identities given in the appendix, we get:
${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)\right.$
$\left.+A_{k}(i)(\mu_{k}\overline{u}^{\top}_{k}(i)+\Sigma_{k}K_{k}^{\top}(i))B^{\top}_{k}(i)+A_{k}(i)\mu_{k}c^{\top}_{k}\right.$
$\left.+B_{k}(i)(\mu_{k}\overline{u}^{\top}_{k}(i)+\Sigma_{k}K^{\top}_{k}(i))^{\top}A_{k}^{\top}(i)+B_{k}(i)\overline{u}^{\top}_{k}(i)c^{\top}_{k}\right.$
$\left.+B_{k}(i)(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i)\Sigma_{k}K_{k}^{\top}(i))B_{k}^{\top}(i)+c_{k}c_{k}^{\top}\right.$
$\left.+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+G_{k}(i)G_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)\mu_{k}\mu_{k}^{\top}A^{\top}_{k}(i)\}\right]\Bigr{)}-\mu_{k+1}\mu_{k+1}^{\top}$ (7)
Remark 1.
eq. 7 shows that the mean-covariance separation of single-mode dynamics ([10]) does not extend to MJLS.
We define a mean trajectory $\tau$ as a finite sequence of $(N+1)$-tuples $\{\mu_{k},\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in{\mathbb{Z}_{0:T-1}}}$ which satisfy eq. 5. We now define our optimal control problem in terms of $\mu_{k}$ and $\Sigma_{k}$, which approximates Problem 1.
Problem 2.
Solve:
$\min_{\mu_{k},\overline{u}_{k}(i),K_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (8)
$s.t.\;\forall k\in\mathbb{Z}_{0:T-1},$
[…]
We write $J_{k}$ in terms of $\mu_{k}$ and $\Sigma_{k}$ as:
$J_{k}$ $=\sum_{i\in\Omega}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
Using the identities derived in appendix A, we can rewrite $J_{k}$ as:
$J_{k}$ $=\sum_{i\in\Omega}\rho_{k}(i)\left[\mathrm{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k})+\mathrm{Tr}(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right.$ (9)
$\left.\hskip 28.45274pt+\mathrm{Tr}({\Sigma}_{k}Q_{k})+\mathrm{Tr}(K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)R_{k})\right]$
The nonlinearity in eq. 7 make Problem 2 results in a non-convex optimization landscape, while the chance constraints eq. 2d are not yet represented in closed form. We tackle these challenges in the following sections by first proposing a tight convex reformulation of Problem 2 without chance constraints, and then showing that these relaxations remain tight with a deterministic formulation of chance constraints too.
III Unconstrained MJLS Covariance Steering
III-A Optimal Control Problem
The unconstrained optimal control problem is given by:
Problem 3.
Solve:
$\min_{\mu_{k},\overline{u}_{k}(i),K_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (10)
[…]
$\;\;\mu_{0}=\overline{\mu}_{0}\;\;\mu_{T}=\mu_{f}$ (11)
III-B Convex Relaxation
To solve Problem 3 in a computationally tractable manner, we propose a two-step optimization process, which optimizes the mean and covariance variables sequentially using convex relaxations. We first introduce the following change of variables to convexify eq. 7:
$L_{k}(i)$ $=K_{k}(i){\Sigma}_{k},\;$ $Y_{k}(i)$ $=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$ (12)
Rearranging eq. 7 using the above-mentioned terms,
${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)+G_{k}(i)G_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)L^{\top}_{k}(i)B_{k}^{\top}(i)+B_{k}(i)L_{k}(i)A_{k}^{\top}(i)+B_{k}(i)Y_{k}(i)B_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)\mu_{k}\overline{u}^{\top}_{k}(i)B^{\top}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)\mu^{\top}_{k}A_{k}^{\top}(i)+A_{k}(i)\mu_{k}c_{k}^{\top}\right.$
$\left.+B_{k}(i)\overline{u}_{k}(i)c^{\top}_{k}+c_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+B_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)\right.$
$\left.+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}c_{k}^{\top}+A_{k}(i)\mu_{k}\mu_{k}^{\top}A^{\top}_{k}(i)]\}\right]\Bigr{)}-\mu_{k+1}\mu_{k+1}^{\top}$
For a given mean trajectory $\tau$, we define covariance trajectory $\Xi$ as a finite sequence of $(2N+1)$-tuples $\{\Sigma_{k},L_{k}(1)\ldots L_{k}(N),Y_{k}(1)\ldots Y_{k}(N)\}_{k\in\mathbb{Z}_{0}^{T-1}}$ which satisfy section III-B. Thus, section III-B can be expressed as
$H_{k}^{\Sigma}=S_{k+1}(j)-F_{\tau}(\Xi)=0$ (14)
where $F_{\tau}(\Xi)$ is a linear operator in $\Xi$ for a given $\tau$. We relax the non-convex definition of $Y_{k}(i)$ in eq. 12 as:
$C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)\preceq 0$ (15)
which can be recast as a LMI using Schur’s complement:
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}\succeq 0$ (16)
We use these new variables in the objective function as follows:
$J$ $=J(\tau)+J(\Xi)$ (17)
where
$J(\tau)=\sum_{k=0}^{T-1}$ $\sum_{i\in\Omega}\rho_{k}(i)\left[\text{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k}+\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right]$ (18a)
$J(\Xi)$ $=\sum_{k=0}^{T-1}\sum_{i\in\Omega}\rho_{k}(i)\left[\text{Tr}({\Sigma}_{k}Q_{k}+Y_{k}(i)R_{k})\right]$ (18b)
III-C Two-step Convex Optimization Framework
We now define the two subproblems which are solved sequentially. Therefore, the two-step optimization process first solves Problem 4 and then uses its solution to solve Problem 5.
Problem 4.
Solve:
$\min_{\mu_{k},\overline{u}_{k}(i)}J(\tau)$ (19a)
[…]
$\mu_{0}=\overline{\mu}_{0},\;\mu_{T}=\mu_{F}$ (19c)
Problem 5.
Given a mean trajectory $\tau$, solve:
$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i)}J(\Xi)$ (20a)
[…]
$\Sigma_{0}=\overline{\Sigma}_{0},\quad\Sigma_{T}\preceq\Sigma_{f}$ (20c)
Assumption 3.
Slater’s condition is satisfied for Problem 5, i.e. there exists a strictly feasible solution for Problem 5.
Lemma 1 (Lemma 1 in [11]).
Let $A$ and $B$ be $n\times n$ symmetric matrices with $A\succeq 0,B\preceq 0$, and $\mathrm{Tr}(AB)=0$. If $B$ has at least one nonzero eigenvalue, then, $A$ is singular.
Theorem 1.
Under Assumptions 1, 2 and 3, the optimal solution to Problem 5 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$, hence, lossless.
Proof.
For Problem 5, we can define the Lagrangian as:
$\mathcal{L}(\cdot)=J(\Xi)+$ $\sum_{k=0}^{T-1}\left[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))\right\}\right]$ (21)
$+\mathrm{Tr}(\lambda^{\top}_{F}(\Sigma_{T}-\Sigma_{f}))$
where $C_{k}(i)$ is defined in eq. 15, $\Lambda_{k}$, $M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively. $\Lambda_{k}$ is symmetric because of symmetry in eq. 14 and $M_{k}$ and $\lambda_{F}$ are symmetric by definition. The relevant first-order KKT conditions for optimality are:
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}$ $=2\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}=0$ (22a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}$ $=\rho_{k}(i)R_{k}-M_{k}(i)+\sum_{j\in\Omega}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (22b)
$\mathrm{Tr}($ $M_{k}^{\top}(i)C_{k}(i))=0,\quad C_{k}\preceq 0,\quad M_{k}\succeq 0$ (22c)
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies the corresponding Lagrange multiplier $M_{k}(i)$ is singular from Lemma 1 and eq. 22c. Rearranging eq. 22a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$. Using this and the substitutions defined earlier in eq. 12, we simplify eq. 22b as: $\rho_{k}(i)R_{k}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Taking the determinant of both sides, we get the following:
$\mathrm{det}(\rho_{k}(i)R_{k})=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
which contradicts $\rho_{k}(i)R_{k}\succ 0$ from Assumption 1. Therefore, we have $C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)=0,\;\forall k,i.$
Remark 2.
Similar relaxations cannot be performed for mean quadratic terms in the eq. 7 because they involve rank-1 matrices which doesn’t allow the use of Schur’s complement to relax the non-convex substitutions as LMIs, motivating our two-step sequential mean-covariance optimization approach.
IV Chance-Constrained MJLS Covariance Steering
We now consider chance constraints for both state and control. For state, we consider intersection of hyperplane constraints and tube-like constraints around the mean trajectory. Mathematically, they’re expressed as:
$P(a^{\top}_{j}x_{k}+b_{j}\leq 0,\forall j)$ $\geq 1-\delta_{x}$ (23a)
$P(\|x_{k}-\mu_{k}\|\leq d_{\max})$ $\geq 1-\epsilon_{x}$ (23b)
Likewise, for control, we consider
$P(p^{\top}_{j}u_{k}(i)+q_{j}\leq 0,\forall j)$ $\geq 1-\delta_{u}(i)$ (24a)
$P(\|u_{k}(i)\|\leq u_{\max}(i))$ $\geq 1-\epsilon_{u}(i)$ (24b)
Previous works [24, 8] have focused on deterministic and convex reformulations of eqs. 23 and 24, but they usually require Gaussian dynamics. MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) In this work, we approximate the MJLS state by a single Gaussian distribution. To achieve this, we seek to minimize the Kullback-Leibler divergence [25] between the MJLS distribution and the approximating Gaussian. It has been shown previously [26] that this is achieved by performing moment matching, i.e. setting the first two moments of the approximating Gaussian to be the same as the original distribution which is being approximated. We leverage this useful property to approximate the state and control of MJLS using Gaussian distributions. We approximate $x_{k}$ by $\mathcal{N}(\mu_{k},\Sigma_{k})$, using moment matching to the approximating Gaussian. Similarly, for control in mode $i$, $\mathbb{E}[u_{k}(i)]=\overline{u}_{k}(i)$ and $\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}]=K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)=Y_{k}(i)$, therefore $u_{k}(i)$ can be approximated by $\mathcal{N}(\overline{u}_{k}(i),Y_{k}(i))$. We now discuss the deterministic formulation of eqs. 23 and 24.
IV-A Deterministic Formulation
Lemma 2 (Lemma 2 in [8]).
Let ${\xi}\sim\mathcal{N}(\bar{{\xi}},P_{\xi})\in\mathbb{R}^{n_{\xi}}$ and $\epsilon\in(0,0.5)$. Then, $\mathbb{P}[a_{j}^{\top}{\xi}+b_{j}\leq 0,\;\forall j]\geq 1-\epsilon$ is implied by
$a_{j}^{\top}\bar{{\xi}}+b_{j}+m_{\mathcal{N}}(\epsilon_{j})\sqrt{a_{j}^{\top}P_{\xi}a_{j}}\leq 0,\;\forall j,\quad\sum_{j}\epsilon_{j}\leq\epsilon$ (25)
where $m_{\mathcal{N}}(\epsilon)=Q_{X\sim\mathcal{N}}(1-\epsilon)$ denotes the quantile function of the standard normal distribution at probability $1-\epsilon$.
Lemma 3 (Lemma 3 in [8]).
Let ${\xi}\sim\mathcal{N}(\bar{{\xi}},P_{\xi})\in\mathbb{R}^{n_{\xi}}$, $\epsilon\in(0,1)$, and $\gamma>0$. Then, $\mathbb{P}[\|{\xi}\|\leq\gamma]\geq 1-\epsilon$ is implied by
$\|\bar{{\xi}}\|+m_{\chi^{2}}(\epsilon,n_{\xi})\sqrt{\lambda_{\max}(P_{\xi})}\leq\gamma$ (26)
$m_{\chi^{2}}(\epsilon,n_{\xi})=\sqrt{Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)}$. Here, $Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)$ denotes the quantile function of the chi-squared distribution with $n_{\xi}$ degrees of freedom, evaluated at probability $(1-\epsilon)$.
Note that while eqs. 25 and 26 are linear in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
$m^{2}_{\mathcal{N}}(\epsilon_{j}){a^{\top}_{j}P_{\xi}a_{j}}\leq(a_{j}^{\top}\bar{\xi}+b_{j})^{2}$ (27)
$m^{2}_{\chi^{2}}(\epsilon,n_{\xi}){\lambda_{\max}(P_{\xi})}\leq(\gamma-\|\bar{\xi}\|)^{2}$ (28)
which are convex in the covariance $P_{\xi}$.
IV-B Iterative Convex Control Synthesis Framework
Using the deterministic formulation of chance constraints in Lemmas 2 and 3, we can now incorporate them in our two-step optimization framework. Note that the chance constraints introduce two-way coupling in the problem, while previously only the covariance problem depended on the mean trajectory and not vice-versa. To solve this, we propose a hierarchical optimization framework which iteratively solves the mean and covariance optimization problem by relaxing the chance constraints. We relax the constraints by adding a slack variable, and then penalize its 1-norm in the objective function. Using the 1-norm promotes sparsity of the slack variable, and has demonstrated favorable convergence properties in other hierarchical optimization frameworks [27, 28]. We now define the relaxed mean and covariance problems as follows:
Problem 6.
Given a covariance trajectory $\Xi$, solve:
$\min_{\mu_{k},\overline{u}_{k}(i)}J_{\mu}^{c}$ (29a)
[…]
$a_{j}^{\top}\mu_{k}$ $+b_{j}+m_{\mathcal{N}}(\delta_{x,j})\sqrt{a_{j}^{\top}\Sigma_{k}a_{j}}\leq\beta_{1,k}$ (29c)
$p_{j}^{\top}\overline{u}_{k}(i)$ $+q_{j}+m_{\mathcal{N}}(\delta_{u,j}(i))\sqrt{p_{j}^{\top}Y_{k}(i)p_{j}}\leq\beta_{2,k}$ (29d)
$\|\overline{u}_{k}(i)\|+$ $m_{\chi^{2}}(\epsilon_{u}(i),n_{u})\sqrt{\lambda_{\max}(Y_{k}(i))}\leq u_{\max}(i)+\beta_{3,k}$ (29e)
where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $J_{\mu}^{c}=J(\tau)+\sum_{m}\alpha_{m}\|\beta_{m}\|_{1}$, with $J(\tau)$ defined in eq. 18a. $\beta_{m,k}>0$ is the slack variable for $m$-th constraint at time step $k$ and $\beta_{m}=[\beta_{m,0}\ldots\beta_{m,T-1}]^{\top}$. $\alpha_{m}>0$ is the weight corresponding to the slack variable vector 1-norm $\|\beta_{m}\|_{1}$ in $J_{\mu}^{c}$.
Problem 7.
Given a mean trajectory $\tau$, solve:
$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i)}J^{c}_{\Sigma}\hskip 56.9055pt$ (30a)
[…]
$m^{2}_{\mathcal{N}}(\delta_{x,j}){a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $-\zeta_{1,j}\leq 0,$ (30c)
$m^{2}_{\mathcal{N}}(\delta_{u,j}){p_{j}^{\top}Y_{k}(i)p_{j}}-(p_{j}^{\top}\mu_{k}+d_{j})^{2}$ $-\zeta_{2,j}\leq 0,$ (30d)
$m^{2}_{\chi^{2}}(\epsilon_{x},n_{x})\lambda_{\max}(\Sigma_{k})$ $-\zeta_{3,j}\leq 0$ (30e)
$m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $-\zeta_{4,j}\leq 0$ (30f)
where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $u_{m}=u_{\max}-\|\overline{u}_{k}(i)\|$, and $J_{\Sigma}^{c}=J(\Xi)+\sum_{j}\alpha_{m}\|\zeta_{m}\|_{1}$ with $J(\Xi)$ defined in eq. 18b. $\zeta_{m,k}>0$ is the slack variable for $m$-th constraint at time $k$ and $\zeta_{m}=[\zeta_{m,0}\ldots\zeta_{m,T-1}]^{\top}$. $\alpha_{m}>0$ is the weight for slack variable vector 1-norm $\|\zeta_{m}\|_{1}$ in $J_{\Sigma}^{c}$.
Assumption 4.
Slater’s condition is satisfied in Problem 7 for $\zeta_{m}=0$. This ensures the convex relaxation in eq. 16 remains lossless while satisfying chance constraints at $\zeta_{m}=0$.
We now show that the convex relaxation $Y_{k}(i)\succeq L_{k}(i)\Sigma_{k}^{-1}L_{k}^{\top}(i)$ remains lossless in Problem 7.
Theorem 2.
Under Assumptions 1, 2 and 4, the optimal solution to Problem 7 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$.
Proof.
We can define the Lagrangian for Problem 7 as follows:
$\mathcal{L}(\cdot)=$ $J_{\Sigma}^{c}+\sum_{k=0}^{T-1}\Bigl{[}\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\lambda_{x,n}e_{x,n}+\lambda_{x,h}e_{x,h}$ (31)
$+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))+\lambda_{u,n}(i)e_{u,n}(i)\right.$
$\left.+\lambda_{u,h}(i)e_{u,h}(i)\right\}\Bigr{]}+\mathrm{Tr}(\lambda^{\top}_{F}(\Sigma_{T}-\Sigma_{f}))$
where $C_{k}(i)$ is given in eq. 15 and $e_{x,h}$, $e_{u,h}$, $e_{x,n}$, and $e_{u,n}$ represent eqs. 30c, 30d, 30e and 30f respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers. $\Lambda_{k}$,$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1. The relevant first-order KKT conditions for optimality are:
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}=2\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}=0$ (32a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}=\rho_{k}(i)R_{k}-M_{k}(i)+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})v_{k}(i)v_{k}^{\top}(i)$
$+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}+\sum_{j\in\Omega}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (32b)
$\mathrm{Tr}(M_{k}^{\top}(i)C_{k}(i))=0,\quad C_{k}\preceq 0,\quad M_{k}\succeq 0,$
$\hskip 56.9055pt\lambda_{x,n},\lambda_{x,h},\lambda_{u,h},\lambda_{u,n}\geq 0$ (32c)
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$ and $v_{k}(i)$ is the eigenvector of $Y_{k}(i)$ corresponding to its maximum eigenvalue. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies $M_{k}(i)$ is singular from Lemma 1 and eq. 32c. Rearranging eq. 32a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$. Using this and the substitutions defined in eq. 12, we simplify eq. 32b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Representing the LHS as $\Delta_{k}(i)$ and taking the determinant of both sides, we get:
$\mathrm{det}(\Delta_{k}(i))=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
From Assumption 1, we have that $\rho_{k}(i)R_{k}\succ 0$. Now we know $\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}\geq 0$ which implies $\Delta_{k}(i)>0$, resulting in a contradiction. Hence, $C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)=0\;\forall k\in\mathbb{Z}_{0}^{T-1},i\in\Omega$. ∎
IV-C Sequential Convex Optimization Framework
We now describe the sequential convex optimization framework which is used to design optimal control laws for MJLS covariance steering with chance constraints. The framework requires a mean trajectory to be initialized (which is obtained by optimizing the unconstrained mean trajectory in Problem 4) and then solve Problem 7 and Problem 6 iteratively with increasing weights on the slack variables until they reach negligible values. This framework is described in Algorithm 1.
Algorithm 1 MJLS Chance-constrained covariance steering
1:Set tol, initial weights $\{\alpha_{m}\}$, and $\eta>1$;
2:Initialize mean trajectory $\tau$ by solving Problem 4;
3:while NotConverged do
4:     Solve Problem 7 with last stored mean trajectory $\tau$ to get new covariance trajectory $\Xi$;
5:     Compute new mean trajectory $\tau$ by solving Problem 6 using revised covariance trajectory $\Xi$ from previous step ;
6:     Store values of slack variables vector $\beta_{m}$ and $\zeta_{m}$;
7:     if $\max_{m}\|\beta_{m}\|_{\infty}$$>$tol or $\max_{m}\|\zeta_{m}\|_{\infty}$ $>$tol then
8:         increase weights $\{\alpha_{m}\}$ by a factor $\eta$;
9:     else
10:         Solution converged. break;
11:return $\tau$ and $\Xi$
V Numerical Simulations
We use the following dynamics:
$A_{k}(1)$ $=\begin{bmatrix}-0.2&1\\ -0.1&0.1\end{bmatrix}$ $B_{k}(1)$ $=\begin{bmatrix}1&0\\ 2&0\end{bmatrix}$ $G_{k}(1)$ $=\begin{bmatrix}0.1&0\\ 0&0.1\end{bmatrix}$
$A_{k}(2)$ $=\begin{bmatrix}0.2&0.1\\ -0.5&0.1\end{bmatrix}$ $B_{k}(2)$ $=\begin{bmatrix}0&1\\ -1&2\end{bmatrix}$ $G_{k}(2)$ $=\begin{bmatrix}0.5&0\\ 0&0.5\end{bmatrix}$
$\mathbb{P}$ $=\begin{bmatrix}0.7&0.3\\ 0.9&0.1\end{bmatrix}$ $c_{k}$ $=\begin{bmatrix}0.01\\ 0.01\end{bmatrix}$ $T$ $=6$
The initial and final conditions are set as $\rho_{0}=[0.5,0.5]$, $\overline{\mu}_{0}=[40,10]^{\top}$, $\overline{\Sigma}_{0}=6I$ and $\mu_{f}=[5,10]^{\top}$, $\Sigma_{f}=2.5I$. We set a half-space state chance constraint of $P(a^{\top}x>b)\geq 0.95$, where $a=[1,0]^{\top}$ and $b=-10$ and a control norm bound constraint $P(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=5$.
We use the CVXPY package[29] with MOSEK solver[30] to run Algorithm 1 on a 2023 Macbook Pro M1 Pro laptop with 16 GB of RAM. All initial weights are set at $\alpha_{m}=1e2$, with $\eta=5$ and $tol=1e-6$. It takes 3.43 seconds and $3$ iterations for Algorithm 1 to converge to the preset tolerance values. Figures 1(a) and 1(b) plot 1000 Monte Carlo trajectories for this problem without and with state chance constraints, respectively. Both figures use the control norm chance constraint. The initial covariance, terminal covariance, and the mean trajectory are shown in green, red, and blue, respectively, while Monte Carlo trajectories are shown in grey. The unshaded area indicates the feasible region under chance constraints, enforced using the state half-space constraint. Figure 2 shows the control norm history for the Monte Carlo simulations. None of the samples disobeyed the state chance constraints, while the control norm exceeded $u_{\max}$ in $1.3\%$ of the total sample size, which meets the prescribed risk bound of $5\%$. These numerical results demonstrate the performance of the controller to steer the covariance within the final target value in the given time horizon, while respecting the chance constraints. Moreover, the state trajectories in Figure 1 demonstrate the multi-modal nature of the problem because of the Markovian jump dynamics and mode-dependent controllers. We also note the number of state distribution modes increasing over time, most apparent in Figure 1(a), clearly demonstrating the exponential Gaussian mixture distribution of MJLS, which is hard to control using conventional stochastic optimal control approaches.
(a) unconstrained case
(b) with chance constraints
Figure 1: State trajectories for Monte Carlo simulations with the state chance constraint $P(x_{1}\geq-10)\geq 0.95$ (shaded region: infeasible.) Both cases use $P(\|u_{k}(i)\|\leq 5)\leq 0.95$.
Figure 2: Control Norm History for the Monte Carlo simulations with state chance constraint $P(x_{1}\geq-10)\geq 0.95$.
VI Conclusion
In this paper, we consider the optimal covariance steering problem for MJLS under chance constraints. We demonstrate the coupled nature of the problem and decompose the problem into a mean-covariance problem, which can be solved in a two-step convex optimization framework. We propose lossless convex relaxations of the covariance subproblem in both unconstrained and chance-constrained cases and discuss an iterative convex optimization framework to solve the chance-constrained covariance steering problem. Numerical simulations for the covariance steering framework are presented. Future work will analyze convergence behavior of the framework and extensions to semi-markov jump systems.
Appendix A MJLS identities
We show some of the identities which are used in deriving the covariance propagation equation for MJLS.
$\mathbb{E}[x_{k}u_{k}^{\top}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[x_{k}(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))^{\top}\mathds{1}_{r_{k}=i}]$
$=\rho_{k}(i)\left[\mu_{k}\overline{u}^{\top}_{k}(i)+{\Sigma}_{k}K^{\top}_{k}(i)\right]$ (33)
$\mathbb{E}[u_{k}(i)c_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))c_{k}^{\top}\mathds{1}_{r_{k}=i}]$
$=\overline{u}_{k}(i)c_{k}^{\top}\mathbb{E}[\mathds{1}_{r_{k}=i}]+0=\overline{u}_{k}(i)c_{k}^{\top}\rho_{k}(i)$ (34)
$\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))(\overline{u}_{k}(i)+$
$\hskip 142.26378ptK_{k}(i)(x_{k}-\mu_{k}))^{\top}\mathds{1}_{r_{k}=i}]$
$=\mathbb{E}[\mathds{1}_{r_{k}=i}]\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+\overline{u}_{k}(i)(\mathbb{E}[x^{\top}_{k}\mathds{1}_{r_{k}=i}]-$
$\mu^{\top}_{k}\mathbb{E}[\mathds{1}_{r_{k}=i}])K_{k}^{\top}(i)+K_{k}(i)(\mathbb{E}[x_{k}\mathds{1}_{r_{k}=i}]-\mu_{k}\mathbb{E}[\mathds{1}_{r_{k}=i}])\overline{u}_{k}^{\top}(i)$
$\hskip 42.67912pt+K_{k}(i)\mathbb{E}[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
$=\rho_{k}(i)\left[\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i){\Sigma}_{k}K_{k}^{\top}(i)\right]$ (35)
Appendix B Discussion on independence of $x_{k}$ and $r_{k}$
In this section, we provide numerical results showing that the relationship between random variables $x_{k}$ and $r_{k}$ is neglible and they can therefore, be considered independent for all practical purposes.
We first compare $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$. A close agreement between the two would indicate our initial assumption. We consider a scalar uncontrolled ($B_{k}(i)=0$) MJLS system without noise or bias ($G_{k}(i)=c_{k}=0$) to simplify the analysis. The system parameters are given below:
$A_{k}(1)=0.8$ $\quad A_{k}(2)=1.2$ $\mathbb{P}=\begin{bmatrix}0.7&0.3\\ 0.9&0.1\end{bmatrix}$ $\quad\rho_{k}(i)=[0.7;0.3]$
Figure 3 shows three plots which compare the sample mean $\mu_{k}$ multiplied by $\rho_{k}(i)$ to $q_{k}(i)$ which is calculated analytically using eq. 4. The plots show high agreement between $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$ indicating that the dependence between $\mathds{1}_{r_{k}=i}$ and $x_{k}$ is neglible.
Figure 3: Figure shows comparison between $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$, where $\mu_{k}$ is calculated using samples and $q_{k}(i)$ is calculated analytically for both modes. The right most plot shows the comparison of sampled and analytical mean $\mu_{k}$, demonstrating that the number of samples considered are sufficient for statistical convergence.
To further demonstrate this, and also provide numerical evidence to prove our mean and covariance propagation equations are correct, we consider the following time-varying system:
$A_{1}(k)$ $=\begin{bmatrix}-0.5&1\\ -1&\sin(0.2\pi k)\end{bmatrix}$ $A_{2}(k)$ $=\begin{bmatrix}\cos(0.2\pi k)&0.5\\ 0.5&0.5\end{bmatrix}$ (36)
$B_{1}(k)$ $=\begin{bmatrix}1\\ 0\end{bmatrix}$ $B_{2}(k)$ $=\begin{bmatrix}0.5\\ 1\end{bmatrix}$
$G_{1}(k)$ $=\begin{bmatrix}1\\ 1.2\end{bmatrix}$ $G_{2}(k)$ $=\begin{bmatrix}0.2\\ 0.3\sin(0.2\pi k)\end{bmatrix}$
$T$ $=10$ $\mathbb{P}$ $=\begin{bmatrix}0.6&0.4\\ 0.5&0.5\end{bmatrix}$
We use a randomized open loop control law sampled uniformly from $[-10,10]$. The noise is $w_{k}\sim\mathcal{N}(1,1000)$ and the initial state distribution is given by $x_{0}\sim\mathcal{N}(\begin{bmatrix}20,100\end{bmatrix}^{\top},10I_{2})$. Figure 4 plots the comparison between sampled mean and covariance values with theoretical calculations using the mean and covariance propagation equations provided in this work for different number of samples. Our simulations demonstrate the numerical accuracy of the propagation equations, while also validating our assumption about $x_{k}$ and $r_{k}$.
Figure 4: Comparison of sampled mean and covariance values with theoretical values calculated using propagation equations for 10,100 and 1e5 samples.