### v4

# Chance-Constrained Covariance Steering for Discrete-Time Markov Jump Linear Systems

Abstract
In this paper, we solve the chance-constrained covariance steering problem for discrete-time Markov Jump Linear Systems (MJLS) using a convex optimization framework. We derive the analytical expressions for the mean and covariance trajectories of time-varying discrete-time MJLS and show that they cannot be separated even without chance constraints, unlike the single-mode dynamics case. To solve the covariance steering problem, we propose a two-step convex optimization framework, which optimizes the mean and covariance subproblems sequentially. Further, we incorporate chance constraints and propose an iterative optimization framework to solve the chance-constrained covariance steering problem for MJLS. Both problems are originally nonconvex, and we derive convex relaxations which are proved to be lossless at optimality using the Karush–Kuhn–Tucker (KKT) conditions. Numerical simulations demonstrate the proposed method by achieving target covariances while respecting chance constraints under additive noise, bias, and Markovian jump dynamics.
I Introduction
Covariance steering is an important problem of designing a control policy that steers a system from an initial given distribution to a target distribution while minimizing a cost function. It was first proposed in [1] as an infinite horizon covariance control problem, and has attracted significant interest recently as a finite horizon covariance steering problem[2, 3, 4, 5].
The covariance steering formulation was previously applied to vehicle path planning [6] and spacecraft trajectory optimization[7, 8, 9]. Such applications typically require path constraints due to safety considerations and physical limitations and have previously been incorporated in the form of probabilistic chance constraints[10]. [11] proposed a lossless convex relaxation of the chance-constrained covariance steering problem, which was later shown [12] to outperform existing methods [4, 6] in computation time.
Despite recent remarkable advances in covariance steering theory for smooth dynamics, covariance steering for hybrid systems has been significantly underexplored. [13] studies covariance steering for a class of hybrid systems that involve state-triggered switching. However, such approaches cannot extend to tackle covariance steering for Markov jump linear systems (MJLS). MJLS is a subclass of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain. They have been extensively studied in the literature with applications in fault-tolerant control [14], network control [15], and macroeconomic models [16]. [17] is an excellent reference on the classical results for discrete-time MJLS.
Constrained control of MJLS has been an active area of research, e.g. with state and control polytopic constraints for linear time-invariant[18] and periodic systems[19], and with second-moment constraints[20]. Chance constraints have previously been incorporated for MJLS[21], but they only consider time-invariant systems in a receding horizon framework using probabilistic reachable sets, which might not scale well with higher dimensions or longer time horizons. Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints and boundary conditions in a computationally tractable manner. To the authors’ best knowledge, such formulations have previously not been investigated in the context of MJLS.
In this work, we focus on the optimal covariance steering problem for MJLS with chance constraints. This is motivated by the need for probabilistic motion planning approaches which face abrupt changes in dynamics, such as spacecrafts[22] and robots[20] with sudden actuator failure. The contributions of this work are threefold: (1) Demonstrate the coupling of MJLS mean and covariance dynamics, showing why mean-covariance separation of single-mode systems does not extend to MJLS, (2) Derive a lossless convex relaxation to solve the MJLS covariance steering problem in a computationally efficient manner, and (3) Incorporate chance constraints in the covariance steering framework and propose a sequential convex programming approach to solve the problem.
Notation
$\mathbb{R}$ and $\mathbb{Z}_{a:b}$ denote the set of real numbers and integers from $a$ to $b$, respectively. $I$ denotes the identity matrix of appropriate size. $\mathds{1}_{\mathcal{\omega}}$ denotes the indicator function. $\lambda_{\max}(\cdot)$, $\mathrm{det}(\cdot)$, and $\mathrm{Tr(\cdot)}$ denote the maximum eigenvalue, determinant, and trace of a matrix respectively. $\|\cdot\|$ denotes the $l_{2}$ norm for a vector, while $\|\cdot\|_{p}$ denotes its $l_{p}$ norm.
II Problem Statement
We consider a discrete-time MJLS defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a wide-sense stationary white noise $w_{k}\in\mathbb{R}^{n_{w}}$ and a deterministic bias $c_{k}$, given by:
$x_{k+1}=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+c_{k}+G_{k}(r_{k})w_{k}$ (1)
where $x\in\mathbb{R}^{n_{x}}$ and $u_{k}(i)\in\mathbb{R}^{n_{u}}$ are the state and control at time $k$ respectively. $r_{k}$ denotes the active mode of a homogeneous Markov chain taking values in the set $\mathcal{S}\triangleq\{1\ldots N\}$ with transition probability matrix $\mathcal{P}$ whose (ii,$j$)th entry is $p_{ij}$. For the white noise $w_{k}$, we have $\mathbb{E}[w_{k}]=0$, $\mathbb{E}[w_{k}w^{\top}_{k}]=I_{n_{w}}$, and $\mathbb{E}[w_{j}w^{\top}_{k}]=0\;\forall i,j,k,j\neq k$ by definition. To study the MJLS dynamics, we focus on the mode probability $\rho_{k}(i)$, and the MJLS state mean $\mu_{k}$ and covariance $\Sigma_{k}$. In MJLS literature, it has typically been easier to work with partial expectations such as $\mathbb{E}[x_{k}\mathds{1}_{r_{k}=i}]$ instead of $\mathbb{E}[x_{k}]$ and we define such variables for the mean and covariance below:
$\rho_{k}(i)\triangleq\mathbb{E}\left[\mathds{1}_{\{r_{k}=i\}}\right],$ $q_{k}(i)\triangleq\mathbb{E}\left[x_{k}\mathds{1}_{r_{k}=i}\right],$
$\overline{x}_{k}(i)\triangleq\mathbb{E}[x_{k}|r_{k}=i],$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
$\mu_{k}\triangleq\mathbb{E}[x_{k}],$ $\hskip-34.1433ptS_{k}(i)\triangleq\mathbb{E}[(x-\overline{x}_{k}(i))(x-\overline{x}_{k}(i))^{\top}\mathds{1}_{r_{k}=i}]$
The mode probability distribution vector is given by $\rho_{k}=[\rho_{k}(1)\ldots\rho_{k}(N)]$. For a given initial mode distribution $\rho_{0}$, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathcal{P}^{k}$. Our framework can be extended trivially to time-inhomogeneous MJLS with $\rho_{k}=\rho_{0}\mathcal{P}_{0}\mathcal{P}_{1}\ldots\mathcal{P}_{k-1}$, but we only consider the time-homogeneous case for notational simplicity. The MJLS covariance steering optimal control problem is then defined as:
Problem 1.
Solve:
$\min_{x_{k},u_{k}(i)}\;J$ $=\sum_{k=0}^{T-1}\mathbb{E}\left[x_{k}Q_{k}x_{k}^{\top}+u_{k}(i)R_{k}u_{k}^{\top}(i)\right]$ (2a)
$\mathrm{s.t}$ $.\;\;\;\forall k\in\mathbb{Z}_{0:T-1},$
$x_{k+1}=A_{k}$ $(r_{k})x_{k}+B_{k}(r_{k})u_{k}+c_{k}+G_{k}(r_{k})w_{k}$ (2b)
$\mathbb{E}[x_{0}]=\mu_{0}$ $,\;\mathrm{cov}(x_{0})={\Sigma}_{0},\;\mathbb{E}[x_{T}]=\mu_{f},\;\mathrm{cov}(x_{T})\preceq\Sigma_{f}$ (2c)
$\mathbb{P}(x_{k}\in\mathcal{X})$ $\geq 1-\epsilon_{x},\quad\mathbb{P}(u_{k}(i)\in\mathcal{U})\geq 1-\epsilon_{u}$ (2d)
where $Q_{k}\succeq 0$, $R_{k}\succ 0$ and $A_{k}(i)$ is invertible $\forall k\in\mathbb{Z}_{0:T},i\in\mathcal{S}$. The distribution for $\rho_{0}$, $w_{k}$ and $x_{0}$ are independent. ${\mu}_{0},\;{\Sigma}_{0}$ are the initial mean and covariance, while $\mu_{f},\Sigma_{f}$ are the final time mean and covariance constraint with $\mathrm{cov}(x)$ denoting the covariance of $x$. $\mathcal{X}$ and $\mathcal{U}$ indicate the feasible sets which must be satisfied with probability $1-\epsilon_{x}$ and $1-\epsilon_{u}$ for state and control respectively.
Assumption 1.
$\rho_{k}(i)>0,\;\forall k\in{\mathbb{Z}_{0:T}},i\in\mathcal{S}$, i.e. the probability of any mode being active is non-zero at all times.
Problem 1 is hard to solve as the decision variables are random variables. To solve it in a computationally tractable manner, we focus on the mean and covariance propagation of the MJLS. However, instead of dealing with $\mu_{k}$ and $\Sigma_{k}$, the MJLS dynamics motivate us to instead work with $q_{k}(i),\overline{x}_{k}(i)$ and $S_{k}(i)$, which are related to $\mu_{k}$ and $\Sigma_{k}$ as:
$\mu_{k}$ $=\sum_{i\in\mathcal{S}}q_{k}(i),\;\;\overline{x}_{k}(i)=\frac{q_{k}(i)}{\rho_{k}(i)}$ (3a)
$\Sigma_{k}$ $=\sum_{i\in\mathcal{S}}\left(S_{k}(i)+\rho_{k}(i)(\overline{x}_{k}(i)-\mu_{k})(\overline{x}_{k}(i)-\mu_{k})^{\top}\right)$ (3b)
We consider a mode-dependent affine feedback control policy:
$u_{k}(i)=\overline{u}_{k}(i)+K_{k}(i)\left(x_{k}-\overline{x}_{k}(i)\right)$} (4)
Assumption 2.
Under eq. 4, $S_{k}(i)\succ 0,\;\forall k\in{\mathbb{Z}_{0:T},i\in\mathcal{S}}$.
Such control laws with a linear feedback law is common in covariance steering literature [11, 12] as it allows us to steer both the mean and covariance simultaneously. We will now use the first and second order moments to solve Problem 1 in a tractable manner. Focusing on the mean, $q_{k}(i)$ propagation can be performed as given below:
$q_{k+1}$ $(j)=\mathbb{E}\left[x_{k+1}\mathds{1}_{\{r_{k+1}=j\}}\right]$
[…]
$\left.\left.\hskip 85.35826pt+c_{k}+G_{k}(i)w_{k}\right)p_{ij}\mathds{1}_{\{r_{k}=i\}}\right]$
$\hskip-14.22636pt={\sum_{i\in\mathcal{S}}p_{ij}\left({A}_{k}(i)q_{k}(i)+\rho_{k}(i)(B_{k}(i)\overline{u}_{k}(i)+c_{k})\right)}$ (5)
Similarly, covariance propagation can be studied using $S_{k}(i)$ as follows:
$S_{k+1}(j)$ $=\mathbb{E}[(x_{k+1}-\overline{x}_{k+1}(j))(x_{k+1}-\overline{x}_{k+1}(j))^{\top}\mathds{1}_{r_{k+1}=j}]$
$\hskip-14.22636pt=\mathbb{E}[x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k+1}=j}]-\rho_{k+1}(j)\overline{x}_{k+1}(j)\overline{x}_{k+1}^{\top}(j)$ (6)
Substituting the identities given in the appendix, we get:
$S_{k+1}(j)=\Bigl(\sum_{i\in\mathcal{S}}\left[p_{ij}\{A_{k}(i)S_{k}(i)A^{\top}_{k}(i)\right$.
$+A_{k}(i)S_{k}(i)K_{k}^{\top}(i)B_{k}^{\top}(i)+B_{k}(i)K_{k}(i)S_{k}^{\top}(i)A_{k}^{\top}(i)$
$+B_{k}(i)(\rho_{k}(i)\overline{u}_{k}(i)\overline{u}_{k}^{\top}(i)+K_{k}(i)S_{k}(i)K_{k}^{\top}(i))B_{k}^{\top}(i)$
$+\rho_{k}(i)(B_{k}(i)\overline{u}_{k}(i)c_{k}^{\top}+B_{k}(i)\overline{u}_{k}(i)\overline{x}_{k}^{\top}A_{k}^{\top}(i)+c_{k}c_{k}^{\top}$
$+c_{k}\overline{u}_{k}^{\top}(i)B_{k}^{\top}(i)+c_{k}\overline{x}_{k}^{\top}(i)A_{k}^{\top}(i)+A_{k}(i)\overline{x}_{k}(i)\overline{u}_{k}^{\top}(i)B_{k}^{\top}(i)$
$+A_{k}(i)\overline{x}_{k}(i)c_{k}^{\top}+A_{k}(i)\overline{x}_{k}(i)\overline{x}_{k}^{\top}(i)A_{k}^{\top}(i)+G_{k}(i)G_{k}^{\top}(i))\}\Bigr)$
$\hskip 71.13188pt-\rho_{k+1}(j)\overline{x}_{k+1}(j)\overline{x}_{k+1}^{\top}(j)$ (7)
Remark 1.
A careful analysis of eq. 7 along with eq. 3b show that the terms involving the MJLS mean information do not cancel out for the covariance propagation, demonstrating that the mean-covariance separation of single-mode dynamics (shown in [10]) does not extend to MJLS.
For a given finite sequence of tuples $\{q_{k}(1)\ldots q_{k}(N)\}_{k\in\mathbb{Z}_{0:T}}$ and $\{\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in\mathbb{Z}_{0:T-1}}$ which satisfy eq. 5, we define a corresponding mean trajectory $\tau$ as a set of two finite sequences of $N+1$ and $N$-tuples $\{\mu_{k},\overline{x}_{k}(1)\ldots\overline{x}_{k}(N)\}_{k\in\mathbb{Z}_{0:{T}}}$ $\{\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in\mathbb{Z}_{0:{T-1}}}$ respectively which satisfy eq. 3a and eq. 5 simultaneously.
We now redefine the cost function $J$ in terms of the newly defined mean and covariance variables. From eq. 2a, we know $J=\sum_{k=0}^{T-1}J_{k}$, where $J_{k}$ is given by
$J_{k}$ $=\sum_{i\in\mathcal{S}}\mathbb{E}\left[(x_{k}^{\top}Q_{k}x_{k}+u^{\top}_{k}(i)R_{k}u_{k}(i))\mathds{1}_{r_{k}=i}\right]$
$=\sum_{i\in\mathcal{S}}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
Using the identities derived in appendices A and A in the appendix, rewrite $J_{k}$ as:
$J_{k}$ $=\sum_{i\in\mathcal{S}}\left[\rho_{k}(i)\{\mathrm{Tr}(\overline{x}_{k}(i)\overline{x}_{k}^{\top}(i)Q_{k})+\mathrm{Tr}(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\}\right$.
$\left.\hskip 14.22636pt+\mathrm{Tr}({S}_{k}(i)Q_{k})+\mathrm{Tr}(K_{k}(i)S_{k}(i)K^{\top}_{k}(i)R_{k})\right]$ (8)
We now define our optimal control problem in terms of the mean and covariance variables using eq. 4, which approximates Problem 1.
Problem 2.
Solve:
$\min_{\mu_{k},q_{k}(i),\overline{x}_{k}(i),\overline{u}_{k}(i),K_{k}(i),S_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (9)
$\mathrm{s.t.}\;\forall k\in\mathbb{Z}_{0:T-1},$
[…]
Problem 2 is still not a tractable problem to solve because of the nonlinear constraint eq. 7 and chance constraints eq. 2d, which are not yet defined as a deterministic constraint in the mean and covariance variables. We tackle these challenges in the following sections by first proposing a lossless convex reformulation of Problem 2 without chance constraints, and then showing that these relaxations remain lossless with the deterministic formulations of chance constraints. Our lossless relaxations convert an equality constraint to an inequality constraint, which is then shown to be tight, i.e. equivalent to the original equality constraint at optimality.
III Unconstrained MJLS Covariance Steering
III-A Optimal Control Problem
The unconstrained optimal control problem is given by:
Problem 3.
Solve Problem 2 without constraints in eq. 2d.
III-B Convex Relaxation
The nonlinear terms in eq. 7 result in a non-convex optimization landscape for Problem 3. We first relax the nonlinear terms involving the covariance variables $K_{k}(i)$ and $S_{k}(i)$ in eq. 7 by introducing the following change of variables:
$L_{k}(i)$ $=K_{k}(i){{S}_{k}(i)},\;$ $Y_{k}(i)$ $=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)$ (10)
Rearranging eq. 7 using the above-mentioned terms,
$S_{k+1}(j)=\Bigl(\sum_{i\in\mathcal{S}}\left[p_{ij}\{A_{k}(i)S_{k}(i)A^{\top}_{k}(i)\right$.
$+A_{k}(i)L_{k}^{\top}(i)B_{k}^{\top}(i)+B_{k}(i)L_{k}(i)A_{k}^{\top}(i)+B_{k}(i)Y_{k}(i)B_{k}^{\top}(i)$
$+\rho_{k}(i)B_{k}(i)\overline{u}_{k}(i)(A_{k}(i)\overline{x}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)+c_{k})^{\top}$
$+\rho_{k}(i)c_{k}(A_{k}(i)\overline{x}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)+c_{k})^{\top}$
$+\rho_{k}(i)A_{k}(i)\overline{x}_{k}(i)(A_{k}(i)\overline{x}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)+c_{k})^{\top}$
$+G_{k}(i)G_{k}^{\top}(i))\}\Bigr)-\rho_{k+1}(j)\overline{x}_{k+1}(j)\overline{x}_{k+1}^{\top}(j)$ (11)
For a given mean trajectory $\tau$, we define covariance trajectory $\Xi$ as a set of two finite sequences of $N+1$ and $2N$-tuples $\{\Sigma_{k}\}_{k\in\mathbb{Z}_{0:T}}$ and $\{L_{k}(1)\ldots L_{k}(N),Y_{k}(1)\ldots Y_{k}(N)\}_{k\in\mathbb{Z}_{0:T-1}}$ which satisfy section III-B. Thus, section III-B can be expressed as
$H_{k}^{\Sigma}=S_{k+1}(j)-F_{\tau}(\Xi)=0$ (12)
where $F_{\tau}(\Xi)$ is a linear operator in $\Xi$ for a given $\tau$. We relax the non-convex definition of $Y_{k}(i)$ in eq. 10 as:
$C_{k}(i)=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)-Y_{k}(i)\preceq 0$ (13)
which can be recast as a LMI using Schur’s complement:
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{{S}_{k}(i)}\end{bmatrix}\succeq 0$ (14)
We note that section III-B still has nonlinear terms involving the mean variables. To solve Problem 3 in a computationally tractable manner, we propose a two-step optimization scheme which solves the mean and covariance problem sequentially. The objective function can be divided into a mean and covariance part using the new variables as follows:
$J$ $=J(\tau)+J(\Xi)$ (15)
where
$J(\tau)=\sum_{k=0}^{T-1}$ $\sum_{i\in\mathcal{S}}\rho_{k}(i)\{\mathrm{Tr}(\overline{x}_{k}(i)\overline{x}_{k}^{\top}(i)Q_{k}+\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\}$ (16a)
$J(\Xi)$ $=\sum_{k=0}^{T-1}\sum_{i\in\mathcal{S}}\left[\text{Tr}({S}_{k}(i)Q_{k}+Y_{k}(i)R_{k})\right]$ (16b)
III-C Two-step Convex Optimization Framework
We now define the two subproblems which comprise our two-step optimization framework. It first solves Problem 4 and then uses its solution to solve Problem 5.
Problem 4.
Solve:
$\min_{q_{k}(i),\overline{x}_{k}(i),\overline{u}_{k}(i)}J(\tau)$ (17a)
[…]
$q_{0}(i)=\rho_{0}(i){\mu}_{0},\;\sum_{i\in\mathcal{S}}q_{T}(i)=\mu_{F}$ (17c)
Problem 5.
Given a mean trajectory $\tau$, solve:
$\hskip-85.35826pt\min_{S_{k}(i),L_{k}(i),Y_{k}(i)}J(\Xi)$ (18a)
[…]
$\hskip-85.35826ptS_{0}(i)=\rho_{0}(i){\Sigma}_{0},\quad\Sigma_{T}\preceq\Sigma_{f}$ (18c)
where $\Sigma_{T}$ is calculated with $S_{T}(i)$ using eq. 3b
Assumption 3.
Slater’s condition is satisfied for Problem 5, i.e. there exists a strictly feasible solution for Problem 5.
Lemma 1 (Lemma 1 in [11]).
Let $A$ and $B$ be $n\times n$ symmetric matrices with $A\succeq 0,B\preceq 0$, and $\mathrm{Tr}(AB)=0$. If $B$ has at least one nonzero eigenvalue, then, $A$ is singular.
Theorem 1.
Under Assumptions 1, 2 and 3, the solution to Problem 5 satisfies $Y_{k}(i)=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)$, hence, lossless.
Proof.
For Problem 5, we can define the Lagrangian as:
$\mathcal{L}(\cdot)=J(\Xi)+$ $\sum_{k=0}^{T-1}\left[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\sum_{i\in\mathcal{S}}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))\right\}\right]$ (19)
$+\mathrm{Tr}(\lambda^{\top}_{F}(\Sigma_{T}-\Sigma_{f}))$
where $C_{k}(i)$ is defined in eq. 13, $\Lambda_{k}$, $M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively. $\Lambda_{k}$ is symmetric because of symmetry in eq. 12 and $M_{k}$ and $\lambda_{F}$ are symmetric by definition. The relevant first-order KKT conditions for optimality are:
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}$ $=2\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){S}_{k}^{-1}(i)=0$ (20a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}$ $=\rho_{k}(i)R_{k}-M_{k}(i)+\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (20b)
$\mathrm{Tr}($ $M_{k}^{\top}(i)C_{k}(i))=0,\quad C_{k}\preceq 0,\quad M_{k}\succeq 0$ (20c)
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies the corresponding Lagrange multiplier $M_{k}(i)$ is singular from Lemma 1 and eq. 20c. Rearranging eq. 20a, we get $\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){S}_{k}^{-1}(i)A_{k}^{-1}(i)$. Using this and the substitutions defined earlier in eq. 10, we simplify eq. 20b as: $\rho_{k}(i)R_{k}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Taking the determinant of both sides, we get the following:
$\mathrm{det}(\rho_{k}(i)R_{k})=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))=0$
which contradicts $\rho_{k}(i)R_{k}\succ 0$ from Assumption 1. Therefore, we have $C_{k}(i)=L_{k}(i)S^{-1}_{k}(i)L_{k}^{\top}(i)-Y_{k}(i)=0,\;\forall k,i.$
Remark 2.
Similar relaxations cannot be performed for mean quadratic terms in the eq. 7 because they involve rank-1 matrices which doesn’t allow the use of Schur’s complement to relax the non-convex substitutions as LMIs, motivating our two-step sequential mean-covariance optimization approach.
IV Chance-Constrained MJLS Covariance Steering
We now consider chance constraints for both state and control. For state, we consider intersection of hyperplane constraints and tube-like constraints around the mean trajectory. Mathematically, they’re expressed as:
$\mathbb{P}(a^{\top}_{j}x_{k}+b_{j}\leq 0,\forall j)$ $\geq 1-\delta_{x}$ (21a)
$\mathbb{P}(\|x_{k}-\mu_{k}\|\leq d_{\max})$ $\geq 1-\epsilon_{x}$ (21b)
Likewise, for control, we consider
$\mathbb{P}(f^{\top}_{j}u_{k}(i)+g_{j}\leq 0,\forall j)$ $\geq 1-\delta_{u}(i)$ (22a)
$\mathbb{P}(\|u_{k}(i)\|\leq u_{\max}(i))$ $\geq 1-\epsilon_{u}(i)$ (22b)
Previous works [23, 8, 10, 6] have focused on deterministic and convex reformulations of eqs. 21 and 22, but they usually assume state distribution to be Gaussian. MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) Therefore, we utilize one-sided Chebyshev/Cantelli’s inequality[24] and the multivariate Chebyshev inequality[25] to convert the chance constraints to deterministic path constraints. Note that while the constraint on the state $x_{k}$ is mode-independent, the control chance constraints are imposed for a given mode $i$. The multi-modal nature of the control variable $u_{k}$ and practical applications motivate us to consider mode-dependent control chance constraints as it would be less conservative than mode-independent control chance constraints. For mode-conditioned control chance constraints, the mean and covariance is given by: $\mathbb{E}[u_{k}(i)|r_{k}=i]=\overline{u}_{k}(i)$ and $\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}|r_{k}=i]=\frac{Y_{k}(i)}{\rho_{k}(i)}$ as shown in the appendix.
IV-A Deterministic Formulation
Lemmas 2 and 3 are useful in the deterministic formulation of half-space and max-norm bound constraints.
Lemma 2.
Let $v\in\mathbb{R}^{n_{v}}$ be a random variable with mean $\overline{v}$ and covariance $V$. Then, $\mathbb{P}[a_{j}^{\top}{v}+b_{j}\leq 0,\forall j]\geq 1-\epsilon$ is implied by:
$a^{\top}_{j}\overline{v}+b_{j}+\sqrt{\frac{1-\epsilon_{j}}{\epsilon_{j}}a^{\top}_{j}Va_{j}}\leq 0,\quad\sum_{j}\epsilon_{j}\leq\epsilon$ (23)
Proof.
See Theorem 3.1 in [26], and Lemma 1 in [8]. ∎
Lemma 3.
Let $v\in\mathbb{R}^{n_{v}}$ be a random variable with mean $\overline{v}$ and covariance $V$. Then $\mathbb{P}[\|{v}\|\leq v_{\max}]\geq 1-\epsilon$ is implied by:
$\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)}\leq v_{\max}$ (24)
Proof.
Given that $\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)}\leq v_{\max}$, we know $\mathbb{P}(\|v\|\leq v_{\max})\geq\mathbb{P}(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)})$…(A). This is described in Figure 1(a). Now using the multivariate Chebyshev inequality [25] for $v$, we get: $\mathbb{P}((v-\overline{v})^{\top}V^{-1}(v-\overline{v})\geq t^{2})\leq\frac{n_{v}}{t^{2}}$…(B). Put $t=\sqrt{\frac{n_{v}}{\epsilon}}$, and considering the complement of the event in (B), we get $\mathbb{P}((v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon})\geq 1-\epsilon$. Note that $(v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon}$ denotes an ellipsoid centered at $\overline{v}$ with shape matrix $V$. We first find the maximum value of $\|v\|$ inside this ellipse. Let $\mathcal{E}=\{v:(v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon}\}$, then: $\max_{v\in\mathcal{E}}\|v\|\leq\max_{v\in\mathcal{E}}\|v-\overline{v}\|+\|\overline{v}\|=\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}+\|\overline{v}\|$. Now construct a ball $\mathcal{B}_{v}$ centered at origin with radius $\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)}$. Geometrically, this is shown in Figure 1(b) for a $n_{v}=2$. Then, $\mathbb{P}(v\in\mathcal{B}_{v})=\mathbb{P}(\|v\|\leq\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}+\|\overline{v}\|)\geq\mathbb{P}(v\in\mathcal{E})$ since $\mathcal{E}\subseteq\mathcal{B}_{v}$. But $\mathbb{P}(v\in\mathcal{E})\geq 1-\epsilon$. Therefore, $\mathbb{P}\left(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}\right)\geq\mathbb{P}(v\in\mathcal{E})\geq 1-\epsilon$. Combining this with (A), we get $\mathbb{P}(\|v\|\leq v_{\max})\geq\mathbb{P}(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)})\geq 1-\epsilon$.
$\overline{v}||+\sqrt{\frac{n}{\epsilon}\lambda_{max}(V)}$$v_{\max}$v||v||p(){p($\cdot$)}
[…]
Figure 1: Schematics for the proof of Lemma 3. Figure (a) shows that $\mathbb{P}(\|v\|\leq v_{\max})\geq\mathbb{P}\left(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n}{\epsilon}\lambda_{\max}(V)}\right)$. Figure (b) shows that $\mathbb{P}(v\in\mathcal{B}_{v})\geq\mathbb{P}(v\in\mathcal{E})$ where the ball $\mathcal{B}_{v}$ is given by $\{v:\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}\}$ and the ellipsoid $\mathcal{E}$ is given by $\{v:(v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon}\}$.
Remark 3.
Note that while eqs. 23 and 24 are linear in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
$\frac{1-\epsilon}{\epsilon}a^{\top}_{j}Va_{j}\leq(a_{j}^{\top}\bar{v}+b_{j})^{2}$ (25)
$\frac{n_{v}}{\epsilon}{\lambda_{\max}(V)}\leq(v_{\max}-\|\bar{v}\|)^{2}$ (26)
which are convex in the covariance $V$.
IV-B Iterative Convex Control Synthesis Framework
Using the deterministic formulation of chance constraints in Lemmas 2 and 3, we can now incorporate them in our two-step optimization framework. Note that the chance constraints introduce two-way coupling in the problem, while previously only the covariance problem depended on the mean trajectory and not vice-versa. To solve this, we propose a hierarchical optimization framework which iteratively solves the mean and covariance optimization problem by relaxing the chance constraints. We relax the constraints by adding a slack variable, and then penalize its 1-norm in the objective function. Using the 1-norm promotes sparsity of the slack variable, and has demonstrated favorable convergence properties in other hierarchical optimization frameworks [27, 28]. We now define the relaxed mean and covariance problems in Problem 6 and Problem 7 respectively as follows:
Problem 6.
Given a covariance trajectory $\Xi$, solve:
$\min_{\mu_{k},\overline{u}_{k}(i)}J_{\mu}^{c}$ (27a)
$\mathrm{s.t.}\;\;\;\forall k\in\mathbb{Z}_{0}^{T-1},i\in\mathcal{S},j$
[…]
$a_{j}^{\top}\mu_{k}$ $+b_{j}+\sqrt{\frac{1-\delta_{x,j}}{\delta_{x,j}}a_{j}^{\top}\Sigma_{k}a_{j}}\leq\beta_{1,k},$ (27c)
$f_{j}^{\top}\overline{u}_{k}(i)$ $+g_{j}+\sqrt{\frac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}p_{j}^{\top}Y_{k}(i)p_{j}}\leq\beta_{2,k}$ (27d)
$\|\overline{u}_{k}(i)\|+$ $\sqrt{\frac{n_{u}(i)}{\epsilon_{u}(i)}\lambda_{\max}(Y_{k}(i))}\leq u_{\max}(i)+\beta_{3,k}$ (27e)
where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $J_{\mu}^{c}=J(\tau)+\sum_{m}\alpha_{m}\|\beta_{m}\|_{1}$, with $J(\tau)$ defined in eq. 16a. $\beta_{m,k}>0$ is the slack variable for $m$-th constraint at time step $k$ and $\beta_{m}=[\beta_{m,0}\ldots\beta_{m,T-1}]^{\top}$. $\alpha_{m}>0$ is the weight corresponding to the slack variable vector 1-norm $\|\beta_{m}\|_{1}$ in $J_{\mu}^{c}$.
Problem 7.
Given a mean trajectory $\tau$, solve:
$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i)}J^{c}_{\Sigma}\hskip 56.9055pt$ (28a)
$\mathrm{s.t.}\;\;\;\forall k\in\mathbb{Z}_{0}^{T-1},\;i\in\mathcal{S},j,\hskip 14.22636pt$
[…]
${\frac{1-\delta_{x,j}}{\delta_{x,j}}a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $\leq\zeta_{1,j},$ (28c)
${\frac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}p_{j}^{\top}Y_{k}(i)p_{j}}-(f_{j}^{\top}\mu_{k}+g_{j})^{2}$ $\leq\zeta_{2,j},$ (28d)
$\frac{n_{x}}{\epsilon_{x}}\lambda_{\max}(\Sigma_{k})$ $\leq\zeta_{3,j}$ (28e)
$\frac{n_{u}(i)}{\epsilon_{u}(i)}\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $\leq\zeta_{4,j}$ (28f)
where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $u_{m}=u_{\max}-\|\overline{u}_{k}(i)\|$, and $J_{\Sigma}^{c}=J(\Xi)+\sum_{j}\alpha_{m}\|\zeta_{m}\|_{1}$ with $J(\Xi)$ defined in eq. 16b. $\zeta_{m,k}>0$ is the slack variable for $m$-th constraint at time $k$ and $\zeta_{m}=[\zeta_{m,0}\ldots\zeta_{m,T-1}]^{\top}$. $\alpha_{m}>0$ is the weight for slack variable vector 1-norm $\|\zeta_{m}\|_{1}$ in $J_{\Sigma}^{c}$.
Assumption 4.
Slater’s condition is satisfied in Problem 7 for $\zeta_{m}=0$. This ensures the convex relaxation in eq. 14 remains lossless while satisfying chance constraints at $\zeta_{m}=0$.
We now show that the convex relaxation $Y_{k}(i)\succeq L_{k}(i)\Sigma_{k}^{-1}L_{k}^{\top}(i)$ remains lossless in Problem 7.
Theorem 2.
Under Assumptions 1, 2 and 4, the optimal solution to Problem 7 satisfies $Y_{k}(i)=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)$.
Proof.
We can define the Lagrangian for Problem 7 as follows:
$\mathcal{L}(\cdot)=$ $J_{\Sigma}^{c}+\sum_{k=0}^{T-1}\Bigl[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\lambda_{x,n}e_{x,n}+\lambda_{x,h}e_{x,h}$ (29)
$+\sum_{i\in\mathcal{S}}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))+\lambda_{u,n}(i)e_{u,n}(i)\right$.
$\left.+\lambda_{u,h}(i)e_{u,h}(i)\right\}\Bigr]+\mathrm{Tr}(\lambda^{\top}_{F}(\Sigma_{T}-\Sigma_{f}))$
where $C_{k}(i)$ is given in eq. 13 and $e_{x,h}$, $e_{u,h}$, $e_{x,n}$, and $e_{u,n}$ represent eqs. 28c, 28d, 28e and 28f respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers. $\Lambda_{k}$$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1. The relevant first-order KKT conditions for optimality are:
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}=2\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i)S_{k}^{-1}(i)$
$\hskip 199.16928pt=0$ (30a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}=\rho_{k}(i)R_{k}-M_{k}(i)+\lambda_{u,n}\tfrac{n_{u}(i)}{\epsilon_{u}(i)}v_{k}(i)v_{k}^{\top}(i)$
$+\lambda_{u,h}\tfrac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}f_{j}f^{\top}_{j}+\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (30b)
$\mathrm{Tr}(M_{k}^{\top}(i)C_{k}(i))=0,\quad C_{k}\preceq 0,\quad M_{k}\succeq 0,$
$\hskip 56.9055pt\lambda_{x,n},\lambda_{x,h},\lambda_{u,h},\lambda_{u,n}\geq 0$ (30c)
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$ and $v_{k}(i)$ is the eigenvector of $Y_{k}(i)$ corresponding to its maximum eigenvalue. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies $M_{k}(i)$ is singular from Lemma 1 and eq. 30c. Rearranging eq. 30a, we get $\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i)S_{k}^{-1}(i)A_{k}^{-1}(i)$. Using this and substitutions defined in eq. 10, we simplify eq. 30b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}\tfrac{n_{u}(i)}{\epsilon_{u}(i)}v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}\tfrac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}f_{j}f^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Representing the LHS as $\Delta_{k}(i)$ and taking the determinant of both sides, we get:
$\mathrm{det}(\Delta_{k}(i))=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))=0$
From Assumption 1, we have that $\rho_{k}(i)R_{k}\succ 0$. Now we know $\lambda_{u,n}\tfrac{n_{u}(i)}{\epsilon_{u}(i)}v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}\tfrac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}f_{j}f^{\top}_{j}\geq 0$ which implies $\Delta_{k}(i)>0$, resulting in a contradiction. Hence, $C_{k}(i)=L_{k}(i){S}^{-1}_{k}(i)L_{k}^{\top}(i)-Y_{k}(i)=0\;\forall k\in\mathbb{Z}_{0}^{T-1},i\in\mathcal{S}$. ∎
IV-C Sequential Convex Optimization Framework
We now describe the sequential convex optimization framework which is used to design optimal control laws for MJLS covariance steering with chance constraints. The framework requires a mean trajectory to be initialized (which is obtained by optimizing the unconstrained mean trajectory in Problem 4) and then solve Problem 7 and Problem 6 iteratively with increasing weights on the slack variables until they reach negligible values. This framework is described in Algorithm 1.
Algorithm 1 MJLS Chance-constrained covariance steering
1:Set tol, initial weights $\{\alpha_{m}\}$, and $\eta>1$;
2:Initialize mean trajectory $\tau$ by solving Problem 4;
3:while NotConverged do
4:  Solve Problem 7 with last stored mean trajectory $\tau$ to get new covariance trajectory $\Xi$;
5:  Compute new mean trajectory $\tau$ by solving Problem 6 using revised covariance trajectory $\Xi$ from previous step ;
6:  Store values of slack variables vector $\beta_{m}$ and $\zeta_{m}$;
7:  if $\max_{m}\|\beta_{m}\|_{\infty}$>>tol or $\max_{m}\|\zeta_{m}\|_{\infty}$ >>tol then
8:   increase weights $\{\alpha_{m}\}$ by a factor $\eta$;
9:  else
10:   Solution converged. break;
11:return $\tau$ and $\Xi$
V Numerical Simulations
We use the following dynamics:
$A_{k}(1)$ $=\begin{bmatrix}-0.2&1\\ -0.1&0.1\end{bmatrix}$ $B_{k}(1)$ $=\begin{bmatrix}1&0.5\\ 2&0\end{bmatrix}$ $G_{k}(1)$ $=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$
$A_{k}(2)$ $=\begin{bmatrix}0.2&0.1\\ -0.5&0.1\end{bmatrix}$ $B_{k}(2)$ $=\begin{bmatrix}0&1\\ -1&2\end{bmatrix}$ $G_{k}(2)$ $=\begin{bmatrix}0.5&0\\ 0&0.5\end{bmatrix}$
$\mathbb{P}$ $=\begin{bmatrix}0.8&0.2\\ 0.9&0.1\end{bmatrix}$ $c_{k}$ $=\begin{bmatrix}0.01\\ 0.01\end{bmatrix}$ $T$ $=6$
The initial and final conditions are set as $\rho_{0}=[0.3,0.7]$, ${\mu}_{0}=[25,40]^{\top}$, ${\Sigma}_{0}=6I$ and $\mu_{f}=[5,10]^{\top}$, $\Sigma_{f}=3I$. We set a half-space state chance constraint of $\mathbb{P}(a^{\top}x+b<0)\geq 0.95$, where $a=[0,-1]^{\top}$ and b=10b=-10 and a control norm bound constraint $\mathbb{P}(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=8$. We consider a zero-mean gaussian distribution $\mathcal{N}(0,I_{n_{w}})$ for $w_{k}$. For sampling $x_{0}$, we also consider a normal distribution with mean $\mu_{0}$ and covariance $\Sigma_{0}$.
We use the CVXPY package[29] with MOSEK solver[30] to run Algorithm 1 on a 2023 Macbook Pro M1 Pro laptop with 16 GB of RAM. All initial weights are set at $\alpha_{m}=10^{2}$, with $\eta=1.5$ and $tol=10^{-6}$. It takes 3.43 seconds and $7$ iterations for Algorithm 1 to converge to the preset tolerance values. Figures 2(a) and 2(b) plot 2500 Monte Carlo trajectories for this problem without and with state chance constraints, respectively. In both cases, the control norm chance constraint is also imposed. The initial covariance, terminal covariance, and the mean trajectory are shown in green, red, and blue, respectively, while Monte Carlo trajectories are shown in grey. The unshaded area indicates the feasible region under chance constraints. Figure 3 shows the control norm history for the Monte Carlo simulations. The Monte Carlo simulations disobeyed the chance constraint in $0.01\%$ of total sample size, which meets the prescribed risk bound of $5\%$. Figure 4 plots $x_{T}$ for all samples, overlayed with the terminal covariance constraint ellipsoid $\Sigma_{f}$, sample covariance and the predicted covariance $\Sigma_{T}$ from the optimizer. These numerical results demonstrate the performance of the controller to steer the covariance within the final target value in the given time horizon, while respecting the chance constraints. Moreover, the state trajectories in Figure 2 demonstrate the multi-modal nature of the problem because of the Markovian jump dynamics and mode-dependent controllers. We also note the number of state distribution modes increasing over time, most apparent in Figure 2(a), clearly demonstrating the exponential Gaussian mixture distribution of MJLS, which is hard to control using conventional stochastic optimal control approaches.
(a) unconstrained case
(b) with chance constraints
Figure 2: State trajectories for Monte Carlo simulations with the state chance constraint $\mathbb{P}(x[2]\geq-10)\geq 0.95$ (shaded region: infeasible.) Both cases use $P(\|u_{k}(i)\|\leq 8)\leq 0.95$.
Figure 3: Control Norm History for the Monte Carlo simulations with state chance constraint $P(x_{2}\geq-10)\geq 0.95$. $u_{pred}(i)$ corresponds to $\|\overline{u}_{k}(i)\|+\sqrt{\tfrac{n_{u}(i)}{\epsilon_{u}}\lambda_{\max}(Y_{k}(i))}$
Figure 4: Sample points from Monte Carlo simulations at final time step, with sampled covariance covariance shown in blue, predicted covariance from optimizer in shaded green, and terminal covariance constraint shown in black. The covariance ellipses are scaled to contain $95\%$ of data of a gaussian distribution with same mean and covariance.
VI Conclusion
In this paper, we consider the optimal covariance steering problem for MJLS under chance constraints. We demonstrate the coupled nature of the problem and decompose the problem into a mean-covariance problem, which can be solved in a two-step convex optimization framework. We propose lossless convex relaxations of the covariance subproblem in both unconstrained and chance-constrained cases and develop an iterative convex optimization framework to solve the chance-constrained covariance steering problem. Numerical simulations for the covariance steering framework are presented, which demonstrate the controller’s performance in steering the covariance of a MJLS to within a given terminal covariance bound while obeying chance constraints with additive noise and bias.
Appendix A MJLS identities
We show some of the identities which are used in deriving the covariance propagation equation for MJLS.
$\mathbb{E}[u_{k}(i)|r_{k}=i]=\frac{\mathbb{E}[u_{k}(i)\mathds{1}_{k}=i]}{\mathbb{E}[\mathds{1}_{r_{k}=i}]}=\frac{\rho_{k}(i)\overline{u}_{k}(i)}{\rho_{k}(i)}=\overline{u}_{k}(i)$ (31)
$\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}|r_{k}=i]$
$=\frac{\mathbb{E}[K_{k}(i)S_{k}(i)K^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]}{\mathbb{E}[\mathds{1}_{r_{k}=i}]}=\frac{Y_{k}(i)}{\rho_{k}(i)}$ (32)
$\mathbb{E}[(x_{k}-\overline{x}_{k}(i))\mathds{1}_{r_{k}=i}]=0$ (33)
$\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(x_{k}-\overline{x}_{k}(i)+\overline{x}_{k}(i))(x_{k}-\overline{x}_{k}(i)+$
$\hskip 14.22636pt\overline{x}_{k}(i))^{\top}\mathds{1}_{r_{k}=i}]=S_{k}(i)+0+0+\rho_{k}(i)\overline{x}_{k}(i)\overline{x}^{\top}_{k}(i)$
$=S_{k}(i)+\rho_{k}(i)\overline{x}_{k}(i)\overline{x}^{\top}_{k}(i)$ (34)
$\mathbb{E}[x_{k}u_{k}^{\top}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[x_{k}(\overline{u}_{k}(i)+K_{k}(i)$
$\hskip 142.26378pt(x_{k}-\overline{x}_{k}(i)))^{\top}\mathds{1}_{r_{k}=i}]$
$=\mathbb{E}[((x_{k}-\overline{x}_{k}(i))+\overline{x}_{k}(i))(\overline{u}_{k}(i)+$
$\hskip 85.35826ptK_{k}(i)(x_{k}-\overline{x}_{k}(i)))^{\top}\mathds{1}_{r_{k}=i}]$
$=0+S_{k}(i)K^{\top}_{k}(i)+\rho_{k}(i)\overline{x}_{k}(i)\overline{u}^{\top}_{k}(i)+0$
$=\rho_{k}(i)\overline{x}_{k}(i)\overline{u}^{\top}_{k}(i)+S_{k}(i)K^{\top}_{k}(i)$ (35)
$\mathbb{E}[u_{k}(i)c_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\overline{x}_{k}(i)))c_{k}^{\top}\mathds{1}_{r_{k}=i}]$
$=\overline{u}_{k}(i)c_{k}^{\top}\mathbb{E}[\mathds{1}_{r_{k}=i}]+0=\overline{u}_{k}(i)c_{k}^{\top}\rho_{k}(i)$ (36)
$\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\overline{x}_{k}(i)))(\overline{u}_{k}(i)+$
$\hskip 113.81102ptK_{k}(i)(x_{k}-\overline{x}_{k}(i)))^{\top}\mathds{1}_{r_{k}=i}]$
$=\mathbb{E}[\mathds{1}_{r_{k}=i}]\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+\overline{u}_{k}(i)(\mathbb{E}[x^{\top}_{k}-\overline{x}^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
$\hskip 85.35826pt+K_{k}(i)\mathbb{E}[x_{k}-\overline{x}_{k}(i)\mathds{1}_{r_{k}=i}]\overline{u}_{k}^{\top}(i)$
$\hskip 28.45274pt+K_{k}(i)\mathbb{E}[(x_{k}-\overline{x}_{k}(i))(x_{k}-\overline{x}_{k}(i))^{\top}\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
$=\rho_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+0+0+K_{k}(i)S_{k}(i)K^{\top}_{k}(i)$
$=\rho_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i)S_{k}(i)K_{k}^{\top}(i)$ (37)